### v2

# CADS: A Systematic Literature Review on the Challenges of Abstractive Dialogue Summarization

Abstract
Abstractive dialogue summarization is the task of distilling conversations into informative and concise summaries. Although focused reviews have been conducted on this topic, there is a lack of comprehensive work that details the core challenges of dialogue summarization, unifies the differing understanding of the task, and aligns proposed techniques, datasets, and evaluation metrics with the challenges. This article summarizes the research on Transformer-based abstractive summarization for English dialogues by systematically reviewing 1262 unique research papers published between 2019 and 2024, relying on the Semantic Scholar and DBLP databases. We cover the main challenges present in dialog summarization (i.e., language, structure, comprehension, speaker, salience, and factuality) and link them to corresponding techniques such as graph-based approaches, additional training tasks, and planning strategies, which typically overly rely on BART-based encoder-decoder models. We find that while some challenges, like language, have seen considerable progress, mainly due to training methods, others, such as comprehension, factuality, and salience, remain difficult and hold significant research opportunities. We further investigate how these approaches are typically assessed, covering the datasets for the subdomains of dialogue (e.g., meeting, customer service, and medical), the established automatic metrics (e.g., ROUGE), and common human evaluation approaches for assigning scores and evaluating annotator agreement. We observe that only a few datasets with community interest span across all subdomains. Despite its limitations, the ROUGE metric is the most commonly used, while human evaluation, considered the gold standard, is frequently reported without sufficient detail on inner-annotator agreement and annotation guidelines. Additionally, we discuss the possible implications of the recently explored large language models and conclude that despite a potential shift in relevance and difficulty, our described challenge taxonomy remains relevant.
1 Introduction
Figure 1: Overview of the big six challenges in dialogue summarization, a short description of the challenges, and the estimation of their progress derived from related sub-challenges. Green means mostly mitigated, orange means good progress, and red stands for marked challenges still exist.
Abstractive dialogue summarization, a task within Natural Language Processing (NLP) and text summarization, entails condensing key information from conversations into succinct and coherent summaries (?). This sub-field of text summarization is gaining prominence and is relevant for various real-world scenarios, including customer service (e.g., social media (?) and e-commerce (?)), healthcare (?), daily life (?), meetings (?), and open-domain conversations (e.g., online-chat (?)). The relevance of this task, which offers insights into discussed concerns or issues, is thereby driven by the increase in digital conversations in these scenarios (?). While the common, manual approach to dialogue summarization is considered to be prone to human errors and consumes significant time and effort, automatic dialogue summarization may effectively reduce this overhead on the user side (?).
Diverging from traditional text summarization, which typically focuses on formal and linear content such as news articles (?) or scientific publications (?), dialogue summarization confronts unique challenges (?): conversational text is inherently dynamic, interactive, and non-linear (?), marked by verbosity, repetition, and informality. Salient information is often distributed across multiple speakers, mixed with casual and off-topic remarks (?). Additionally, the use of elliptical and fragmented sentences, characterized by, e.g., incomplete thoughts and context-specific abbreviations, further hinders the summarization process. These complexities make traditional text summarization approaches less effective, which are not trained to bridge the gap between edited text and dialogue and struggle to transfer learned patterns through few-shot learning (?).
Despite advancements achieved by leveraging and adapting general-purpose pre-trained neural language models (?, ?), and the recent trend to explore Large Language Models (LLMs) for this task (?), there remains a notable gap in the field’s foundational understanding of the challenges when processing conversation transcripts. While research papers introduce techniques addressing similar challenges, their definitions and interpretations of these challenges can vary markedly. For instance, while some describe the informal and ungrammatical nature of spoken language as a language challenge (?, ?), others frame the challenge as different individual styles and unstructured expressions (?, ?). Both points of view are thereby valid but, on their own, focus just on a subpart of the language challenge and blur the understanding of it. Current surveys leave out the challenges in dialogue summarization and their variations in definitions, focusing instead on details on datasets, metrics, and techniques. As a consequence, there is no clear understanding of which aspects of dialogue summarization are well-addressed and where significant research potential lies. Our systematic review consolidates the existing knowledge researched in the field considering Transformer-based models, aiming to provide a comprehensive overview of the challenges and their progress. We further consider the center interest in LLMs and discuss their impact on the relevance of the dialogue summarization challenges.
This article focuses on four key areas, namely challenges, dataset overviews, techniques, and evaluation methods. In challenges, we categorize known challenges for Transformer-based models found when dealing with dialogue transcripts into six broader challenge blocks, i.e., language, structure, comprehension, speaker, salience, and factuality. Our proposed taxonomy is displayed in Figure 1, including the six challenge blocks, a short description, and exemplary sub-challenges of the challenge blocks. In datasets overview, our review displays used datasets focused on dialogue summarization, organized by their subdomain, and further an overview of techniques to generate datasets to cope with the existing data scarcity artificially and how to optimize data usage. In techniques of dialogue summarization, we then match 97 techniques proposed since 2019 to the six challenge blocks. Furthermore, in evaluation methods, we give an overview of typical evaluation metrics used, spanning count-based metrics (e.g., ROUGE (?)), model-based (e.g., BERTScore (?)), QA-based (e.g., QuestEval (?)), and human evaluation metrics for performance and annotator agreement. This systematic review contributes to a better understanding of the dialogue summarization problem, unifies the inherent definitions, overviews established techniques, and demonstrates the scarcity of datasets and fitting evaluation metrics.
Organization of the Literature Review. The remainder of this review is structured as follows. We introduce the current state of dialogue summarization in Section 1.1 and show our methodology in Section 2. In Section 3, we provide the general problem definition and our challenges taxonomy consisting of six challenge blocks inherent in conversation scripts, further linking proposed approaches to handle the challenges. Section 4 overviews prominent datasets while the display of evaluation approaches in Section 5 helps readers choose suitable indices to evaluate the effectiveness of a model. Finally, Section 6 discusses future research directions followed by conclusions in Section 7. All resources for our review are publicly available. [^1] https://anonymous.4open.science/r/LitRev-DialogueSum/
1.1 Existing Reviews in Dialogue Summarization
Comprehensive literature reviews on dialogue summarization remain scarce, with notable contributions by ? (?) and ? (?). Existing surveys mainly explore the subdomains, their datasets, techniques, and metrics within dialogue summarization. They provide a high-level overview and categorization of prominent subdomains like meeting, email, chat, medical, and customer service but often neglect niche areas such as e-commerce and debates (?). On the datasets, ? (?) provide a framework linking linguistic dialogue types (?) (e.g., persuasion, negotiation, inquiry) with established datasets, offering insights into dataset suitability. ? (?) analyze dialogue summarization techniques up to early 2022, categorizing techniques by conversational context and component modeling (e.g., speaker, addressee, and utterance modeling to solve the ’Who says what to whom’ paradigm), covering a subset of the conversation challenges indirectly. ? (?) extend this analysis with techniques from 2023, exploring approaches like hierarchical models for long inputs (?), feature injection for enhanced understanding, auxiliary tasks for broader training objectives, and data augmentation from related tasks. Regarding evaluations, works on dialogue summarization leverage automatic metrics from traditional document summarization, with the study by ? (?) discussing their individual strengths and limitations.
While existing surveys provide an understanding of dialogue summarization regarding subdomains, datasets, techniques, and metrics, the field misses a link between these categories through a common framework, such as the challenges inherent in processing dialogues. Our review addresses this gap by introducing a comprehensive taxonomy of dialogue challenges and categorizing recent techniques until 2024 based on these challenges (Section 3). This approach highlights well-explored areas and points out gaps in current research. We analyze datasets (Section 4) and evaluation metrics (Section 5), linking them to the identified challenges. Additionally, we discuss the implications of the shift from smaller encoder-decoder models such as BART (?) to LLMs and the implications for the relevance of the challenges with this new backbone model type (Section 6). Our review concentrates exclusively on dialogue summarization and does not encompass broader areas such as general text summarization (?) or dialogue generation (?).
2 Methodology
We use the PRISMA checklist (?) for organizing our literature review, as it is an established approach to set up a comprehensive systematic literature review while reducing the potential for incomplete data and biases in content selection and presentation (?). This ensures that our approach is structured, transparent, and reproducible.
Figure 2: An overview of the different stages of our methodology and their inherent sub-stages.
Our methodology comprises two stages as detailed in Figure 2.
Retrieval Stage. We retrieve literature automatically and keyword-based from two established academic databases, i.e., Semantic Scholar and DBLP, using specific retrieval criteria and defining keyword queries for systematic searching. The retrieved works are then automatically de-duplicated and saved to a list.
Manual Stage. The second stage involves manually screening full documents using strict inclusion and exclusion criteria. In this stage, we follow additional best practices for manual filtering beyond those in the PRISMA checklist. These include tracking specific items such as problem category, challenges discussed, and proposed technique category (?, ?) and guidelines for human annotators (?, ?).
We evaluate 1262 unique papers from 2019 until 2024, offering an up-to-date perspective on the state of the art in dialogue summarization using Transformer-based architectures. We decide against including references from selected papers that did not appear in our crawl as additional sources to maintain an unbiased selection process and avoid complicating the stopping criterion for literature inclusion. Considering our exclusion criteria, such as multilingualism, non-abstractive, and non-Transformer-based approaches, we select 133 papers for our literature review.
2.1 Stage 1: Retrieving Candidate Documents
Information Sources. Our primary data sources are Semantic Scholar [^2] https://www.semanticscholar.org/ and DBLP [^3] https://dblp.org/ (DataBase systems and Logic Programming), both recognized for extensive coverage in computer science (?, ?). While Semantic Scholar allows for advanced search functionalities and complex queries (?, ?, ?), DBLP is the most comprehensive database for computer science publications, frequently used in other surveys as sole source (?, ?, ?) and encompassing major peer-reviewed journals and conference proceedings. This dual-source strategy ensures thorough retrieval and coverage specific to dialogue summarization, strengthening the certainty and reliability of our evidence base. To ensure our literature review remains up-to-date, we regularly re-crawl databases while conducting the review to incorporate the latest publications into our analysis.  [^4] Last crawled on March 25, 2024. While we acknowledge the potential for these databases to miss the most recent uploads, our strategy minimizes this risk. We exclude databases such as Google Scholar, which overly returned grey literature, and Web of Science, which emphasized extractive summarization techniques during our pre-testing, to maintain a focus on peer-reviewed literature in abstractive dialogue summarization. We design an automatic pipeline to retrieve query-related works and remove duplicates. We limit the number of results per query to the top 100 based on the relevance ranking provided by the platforms, which was rarely required. We recognize that the ranking systems may be biased towards prestigious journals or highly cited authors. Hence, we retrieve all papers published after June 2023, not relying on the ranking systems to account for their limited time to gather citations.
Data Collection Process. Our search strategy uses key terms related to dialogue summarization identified during our initial related work analysis (Section 1.1). Anchor terms such as ’dialogue summarization’ and ’multi-party conversation’ are combined with specific search terms such as ’technique’. We use four anchor terms and 28 search terms, resulting in 112 search queries for each database. We present the complete list and search keywords in our repository [^5] https://anonymous.4open.science/r/LitRev-DialogueSum/ . Employing a Python-based pipeline (Section 8), we systematically extract documents from Semantic Scholar and DBLP APIs, merging and unifying the results into tabular data. Between January 2019 and March 2024, we retrieved 732 publications from DBLP and 968 from Semantic Scholar. After removing 438 duplicates identified across both sources, we have 1262 unique publications at the end of the stage. These results are tagged with their respective queries and compiled into a CSV file for subsequent selection and analysis.
2.2 Stage 2: Manual Filtering
Eligibility Criteria. For our manual filtering process, we define eligibility criteria to ensure the focus of our literature review. We look at long and short, open-access papers published since 2019 that align with the rise of Transformer-based text summarization models such as BART (?) and PEGASUS (?), which have been the core model for most approaches in the field. This timeframe further allows us to concentrate on current challenges rather than including some that may have lost relevance (e.g., negation (?)) because of new advancements in the field. We exclude non-English primary datasets, multi-modal studies, and extractive or non-Transformer-based methods to maintain a focus on English-language research on abstractive dialogue summarization.
Manual Paper Processing. A team of three reviewers  [^6] Each team member discloses potential conflicts of interest to ensure an unbiased review process., consisting of postgraduate students and doctoral candidates with a computer science background, undertake the manual screening process, starting with titles and abstracts to narrow down the automatic selection from 1262 to 186 documents using our eligibility criteria. They then review the full texts, checking if the works still meet the eligibility criteria, and decide whether to include, exclude, or discuss each paper. Disagreements are resolved through consensus. They also track features such as used datasets, employed backbone models, proposed techniques, and evaluation approaches. This step filters out 53 papers, leaving 133 to compose the literature review and detail challenges, techniques, datasets, and evaluation metrics. To further minimize bias, a secondary review is performed as needed, reaffirming initial selections and categorizations, thereby enhancing the study’s integrity.
3 Challenges and Progress
Abstractive dialogue summarization aims to condense multi-party conversations into their key points, ranging from casual chats to expert discussions, and whether these points were explicitly stated or inferred (?). The task was first thoroughly defined in 2002 by ? (?), outlining challenges unique to spoken dialogues such as handling speech disfluency, the lack of clear sentence boundaries, distribution of salient information across various speakers and turns, resolving references, identifying discourse structures, and dealing with inaccuracies caused by speech recognition errors. While some definitions take a mathematical approach (?), viewing dialogues as sequences of turns that are compressed and each associated with a specific speaker, others adopt a broader description of the task as ”extracting useful information from a dialogue” (?). Although all definitions contain relevant, overlapping aspects, they lay down different viewpoints, each considering a different set of challenges to define dialogue summarization. Our goal is to organize these definitions and provide a concise, structured overview of the progress in abstractive dialogue summarization, extending the earlier challenge-based concept (?) and proposing the Challenges of Abstractive Dialogue Summarization (CADS) Taxonomy (Figure 1) tailored to the current Transformer-based model architecture. This taxonomy consists of six challenges: language, structure, comprehension, speaker, salience, and factuality. The first five challenges are thereby related to the input, while the last one concerns reliability. In the following Sections 3.1, 3.2, 3.3, 3.4, 3.5 and 3.6, we detail the CADS Taxonomy as shown in Figure 1.
For this section, we consider 91 papers in total, from which 73 papers define the challenge characteristics (sub-challenges in italic) and 74 describe explored techniques. We exclude papers mainly discussing evaluation strategies and datasets from this subset and cover them in Sections 4 and 5.
3.1 Language
Characteristics. The language challenge describes the idiosyncratic nature of spoken language and individual speech patterns. This includes linguistic subtleties such as informal expressions (e.g., ’yeah,’), ungrammatical structures, colloquialisms, personal vocabulary, and linguistic noise (e.g., filler words) (?, ?, ?, ?, ?). It further covers content repetition (i.e., speakers restate or rephrase information for emphasis or clarity) (?, ?, ?), and domain-specific terminology (e.g., medical terms) (?, ?, ?, ?). These elements demand that models accurately interpret and adapt to the linguistic features of dialogues (?). Figure 3 presents a dialogue snippet showing examples of these three sub-challenges. Failing to handle this challenge may result in a loss of coherence and clarity, repetition of content, and factuality issues.
Figure 3: Example dialogue to demonstrate examples of the idiosyncratic nature of spoken language and individual speech patterns. Red displays disfluencies, blue highlights personal speech patterns, orange displays the colloquialism, green represent informal expressions
Approaches. Transformer-based models often struggle with the language characteristics due to a gap between their pre-training data, typically well-edited texts (e.g., news articles, research papers, Wikipedia entries) (?), and the characteristics of spoken dialogue not reflected in these edited texts (?). The transfer between these text styles is hindered by the scarcity of diverse, dialogue-oriented pre-training data, resulting in a lack of exposure to dialogue data during training, consequently reducing models’ performance. Repetition of already provided information is less of an issue for current language models (?).
Current research aims to bridge this gap between the formal language of the pre-training data and dialogues by retraining models with dialogue-focused tasks, considering both single- and multi-task setups. Explored tasks span masking key dialogue elements (e.g., pronouns entities, high-content tokens, and words) (?) and part-of-speech tagging (?). More recent tasks aim to transform informal, first-person dialogue into a structured, third-person narrative through changing speaker names, adjusting grammar, and adding emotional context (?), or simulating a dialogue by modifying pre-training documents into a conversation structure (?). Teaching models to understand the ’who-did-what’ structure through a specific task is further gaining traction (?). Other explored strategies include pre-processing the input for anaphora resolution (?) and adapted pre-training (?, ?, ?).
3.2 Structure
Characteristics. Aiming at extracting the built-up of a conversation, the structure challenge revolves around segmenting a dialogue transcript and creating an ordered summary based on the extracted structures (?). This challenge involves identifying the topic flow within a conversation (?, ?, ?, ?), tracking the dialogue phases (e.g., problem identification, decision making) (?, ?)), and analyzing the utterance interaction (i.e., relationship and dependencies between utterances) (?, ?, ?). In Figure 4, we show a dialogue snippet exhibiting topic flow, dialogue phases, and utterance interactions. This can result in summaries lacking coherence, completeness, and depth (?, ?, ?, ?, ?).
Figure 4: Excerpt from a conversation to demonstrate the structural challenge, with a detailed view on one topic out of multiple topics touched in the whole conversation. The red arcs mark the connection between individual turns, demonstrating how different dialogue phases (grey text), such as the conclusion, relate to many previous conversation phases.
Approaches. Transformer-based models face two primary hurdles when handling the structural challenge, i.e., adapting to the variability in dialogue structures, such as differing topic sequences and conversation phases depending on the conversation type, and managing long-range dependencies to link utterances. The first difficulty is related to the lack of varied datasets across different dialogue types on which models can train to improve generalization as well as that most pre-trained models used are expecting clearer structural signals due to the more structured data used during pre-training (Section 3.1). The second problem relates to a weakness of transformed-based architectures: their limitation in handling long sequences, which stems from the practical limitation of a fix-sized context window (?), the diluted attention on a single token when many tokens are processed (?), and the bias towards recent information (?).
Recent methods employ graph structures to understand and use dialogue structures more effectively. They combine static and dynamic graphs for a detailed examination of conversation dynamics (?) where the static graphs represent unchanging aspects like speaker relationships, and dynamic graphs track how dialogues evolve. Abstract Meaning Representation (AMR) graphs are employed to capture overarching themes (?), detailed sentence-level connections, and entity interactions, enhancing content comprehension (?). Meanwhile, a hybrid approach combining traditional language models with graph neural networks captures the utterance interactions and addresses the long-range dependency sub-challenge (?). Similarly, ConceptNet (?) is used for thematic linking (?).
To manage long dialogue transcripts beyond context-size limits and the inherent long-distance relations, researchers are moving away from simply truncating inputs to context length to avoid losing vital information. A leading method segments the summarization process into stages (e.g., topics and dialogue phases) for coherent processing (?, ?, ?). The segmentation and multi-pass strategy iteratively summarizes parts of the dialogue and combines partial summaries with a subset of the remaining content in each iteration to build a complete summary (?). The $Summ^{N}$ framework (?), summarizes dialogue segments independently before refining them together, offering a detailed and scalable solution.
Other techniques involve additional training tasks (?, ?, ?, ?), sentence importance measures to group sentences into sub-topics (?, ?), while hierarchical setups to differentiate dialogue parts based on who’s speaking (word level) and what’s being discussed (turn level) are also popular (?).
3.3 Comprehension
Characteristics. The comprehension challenge involves accurately understanding and contextualizing utterances so models can grasp and leverage these insights for summarization (?). This includes direct and implied content, e.g., unmentioned background knowledge, organizational knowledge, and prior discussions not explicitly referred to (examples in Figure 5). To infer and understand this content, the whole context of a dialogue is required, spanning both the local context, focusing on neighboring words and sentences, and the global context, covering the entire dialogue’s informational flow (?).
Figure 5: Example conversation to demonstrate the comprehension challenge including direct and implied content. Directly stated (orange background) are the explicit updates by Bob and Charlie about their respective tasks. Green marks implied content such as a deadline for a demo and a current focus on performance and finalizing features (unmentioned background knowledge), Redis implies an understanding of its role in performance enhancement (organizational knowledge and reference to prior discussions on data requirements).
Approaches. Transformer-based models struggle with comprehension and contextualization due to the underlying dependency parsing challenges, particularly with unclear sentence boundaries as they appear in spoken language, which can lead to catastrophic forgetting and result in instability in handling long-range dependencies (?). Additionally, these models often interpret the text literally, affecting their ability to grasp nuanced meanings (?, ?). As these models often rely on syntactic relationships between words, they miss the nuanced understanding humans use to interpret implied meanings of sarcasm, irony, euphemisms, and similar pragmatic elements (?). Consequently, this can lead to summaries that miss deeper meanings or implications, resulting in content that may be shallow or misleading (?).
Despite its importance, our review reveals limited engagement from the dialogue summarization community to improve the contextualizing capability of models considering both the short and long context. Current efforts primarily rely on Transformer models’ self-attention capabilities to comprehend local and global contexts (?). Despite this reliance, the context-aware extract-generate framework (?) has been proposed to reduce the risk of missing long-range contexts in complex dialogues, which uses context prompts to capture the immediate details and the overall narrative of text spans. It first pinpoints relevant text parts that guide the extraction of key information, which is then used as a prompt base for generating detailed summaries.
3.4 Speaker
Characteristics. Participants, their interactions, and how they influence the dynamics of a conversation are part of the speaker challenge. This includes the individual participant roles (e.g., ”project manager”, ”customer service agent”) (?, ?), topic-dependent role changes (?, ?, ?, ?, ?, ?, ?), their interactions (e.g., contributions depending on viewpoints (?, ?, ?), backgrounds (?, ?, ?)), and references to entities and other participants (?, ?, ?, ?). Examples of the speaker characteristics are shown in Figure 6. Failing to capture these speaker characteristics may lead to missing context, misaligned entity associations, and, ultimately, incorrect factual summaries. The significance of examining speaker dynamics and roles is well-recognized in other fields such as linguistics (?) and is, together with the language challenge of the preceded analysis, a common hook to demonstrate and motivate techniques for dialogue summarization.
Figure 6: Example conversation of the speaker challenge with defined roles (Alex as customer, Jordan as customer support agent, Taylor as technical specialist). The dialogue shows a change of roles when Jordan calls Taylor (marked with orange), shifting from Jordan as a primary troubleshooter to a facilitator for Taylor’s technical interventions. Further, multiple direct references to entities highlighted with blue (e.g., ”device”, ”error codes”) and indirect references to participants in red (e.g., ”you”) are incorporated.
Approaches. The speaker challenge arises from identifying participants, tracking their actions and entity references, and revealing participant relationships. These challenges can be aligned with the difficulties of named entity recognition (NER), coreference resolution, and dependency parsing. Although advancements with Transformer-based models regarding NER, the explored techniques struggle to accurately identify and categorize entities across different text types and domains and with unstructured texts (?, ?, ?), rendering realistic dialogue scenarios as a challenging field. Coreference resolution poses difficulties for techniques in accurately identifying mentions and contextual links (?) and understanding the linguistic structures and relationships (?), extending the understanding challenge (Section 3.3).
Works typically use dense vectors (?, ?) to accurately represent speaker roles in dialogue and act as labels for utterances (?, ?), offering insights into each speaker’s function within the dialogue. Speaker vectors are either randomly initialized and trainable (?, ?) or produced by small neural networks (?, ?) and represent different roles, e.g., ”industrial designer” (?), ”judge” (?), which are appended to the embedding of the speaker’s turn. For capturing speaker dynamics and resolving coreferences, graph-based models are an established approach, e.g., to build a graph representing each speaker’s main ideas and their discourse, alongside their interactions, reflecting inner- and inter-speaker structures (?, ?).
Some other popular techniques involve enhancing Transformers’ self-attention to analyze intra- and inter-speaker dynamics (?, ?), additional training tasks (?, ?, ?), and pre- and post-processing techniques to replace pronouns and correct mistakes (?, ?, ?, ?) for coreference resolution.
3.5 Salience
Characteristics. The salience challenge involves identifying and summarizing the most relevant information within a dialogue (?). This includes identifying content critical for insightful summaries (?), where the understanding of critical shifted lately from content for general audience summaries to summaries aligning with the viewpoints, values, and ethics of specific participant (?), and locating text spans containing relevant content (?). Figure 7 demonstrates the differences between personalized and generally applicable summaries, as the generated texts for Simon and Nadia are tailored to their role, stands, and tasks, whereas the general summary provides a neutral overview. Failure to effectively locate the scattered information or understand what is considered salient can impact the depth and relevance of the generated summaries, focusing on unimportant content while omitting relevant ones (?, ?, ?, ?).
Figure 7: Example of the salience challenge with a conversation being summarized for a general audience and tailored summaries for participants with a specific focus on content relevant to them.
Approaches. The first hurdle for current models to the salience challenge is the request for subjective salience and ranking content’s importance to align with the role and knowledge of the summary addressee. This is difficult as there is no training task for inferring personal attributes and ranking information according to them. Further, locating this content is difficult as salient information is typically scattered across multiple turns within the idiosyncrasies of spoken language (?, ?, ?), which forms an extension of the language challenge (Section 3.1).
To nest the general understanding of what is considered salient into models, established strategies include additional training stages to better distinguish between salient and non-salient content. ? (?) introduce the uncovered loss to point out if salient information is missing and the contrastive loss to distinguish between relevant and irrelevant sentences, respectively. Negative samples, categorized as redundant (i.e., text with unnecessary utterances) and null (i.e., text with relevant utterances removed), help models prioritize important content (?, ?). To better determine salient content for specific participants, works propose a fine-tuning task in which the topic or user perspective to be focused on is passed as an input along the dialogue text (?), apply question-driven pre-training (?) and use dynamic prompts to direct model attention to key dialogue aspects (?). Additional strategies involve direct text manipulation (?), personal named entity planning to concentrate on specific entities (?), adapting attention mechanisms (?), and integrating human feedback (?) or graph structures (?) to further tailor summaries to the dialogue’s essential parts.
3.6 Factuality
In contrast to the previous input-related challenges, the factuality challenge describes the problem of correcting false content in a generated summary caused by a non-robust architecture. Factual incorrect content contains intrinsic errors, such as hallucinating events not present in the original transcript, and extrinsic errors, such as incorrect coreference resolution misrepresenting actual event details, and wrong conclusions stemming from negation (?). Intrinsic and extrinsic errors are shown in a sample dialogue in Figure 8. Given the importance of ensuring dialogue summarizers accurately reflect the conversation, research focuses on creating safety measures to maintain summary credibility despite these challenges.
Figure 8: Example of the factuality challenge where the generated summary exhibits intrinsic errors through hallucinating events (blue background) such as the finalization of a marketing strategy and a detailed review of financial strategies. The extrinsic error reporting the wrong date of the meeting is marked with red.
Approaches. Research interest predominantly targets post-processing approaches that rewrite the generated summary to correct factual errors. Notable approaches here are the usage of LLMs for error detection and correcting (?) and reformulating the summary to closely match the original dialogue’s predictive capabilities for subsequent content (?). Other approaches explore the use of auxiliary tasks (?) to estimate the factual aspects and predict the missing aspects in a summary (?), losses such as encouraging a model to generate sentences about content not yet covered and differentiating factual from non-factual sentences (?, ?), and human feedback (?). Further, architecture-modifying approaches are explored, such as a new encoder to grasp dialogue states (?) or to handle graph structures, where events of dialogue are captured and organized in a graph. A special attention mechanism is used to fill in summary (?), and a hierarchical approach builds on sub-summaries (?). Another noteworthy technique is the generation of a sketch to structure out the final summary along this plan (?). Negation is a long-standing challenge for language models often neglected in computational studies (?, ?).
4 Datasets
In this review, we collected both methods and downstream datasets from the selected publications. We only consider datasets used at least five times in our analysis to avoid including resources with little community impact. This yields 18 datasets that we group into six categories based on the dialogue subdomains they stem from: daily chat, online chat, meeting, screenplay, customer service, and medical. We also create another category for datasets that don’t fit neatly into these subdomain groups.
These datasets typically condense a short dialogue of about 1k tokens between two real persons into a third-person summary of about 100 tokens. More complex scenarios, such as meetings, parliamentary debates, and TV series dialogues, involve around 10k to 20k tokens in the transcript and five or more participants. Due to privacy concerns, a significant portion of currently publicly available datasets are reenactments based on real events, e.g., by Amazon Mechanical Turk workers, or established datasets such as SAMSum (?) and AMI (?) are modified to generate new data, e.g., by changing the dialogue transcript to a third-person report (?). This underscores that the variety of existing datasets is comparably limited, and these datasets may not capture real-life scenarios. The 18 available established datasets are detailed in Sections 4.1, 4.2, 4.3, 4.5, 4.7, 4.6, 4.4 and 4.8. Given the limited number of available datasets, we present methods for creating artificial datasets and strategies for optimizing the use of existing datasets in Section 4.9.
For this section, we consider 93 papers to identify the established datasets, excluding works that do not mention a dataset, datasets that are no longer publicly available, and datasets in a language other than English. We did not assess dataset quality or perform a detailed analysis of the inherent challenges, leaving that for future research. In Figure 9, we connect the datasets with their respective primary challenge, as noted by their creators, to provide an overview of the distribution of challenges in datasets.
Figure 9: Matching between established datasets and their primary reported challenge.
4.1 Daily Chat
DialogSum (?) integrates dialogues from DailyDialog (?), DREAM (?), MuTual (?), and English-speaking practice websites, featuring two-speaker interactions across daily-life scenarios like work, leisure, and travel. Annotators were given guidelines on writing the summaries, such as length constraints (no longer than 20% of conversation length), maintaining an observer perspective, and explicitly describing important emotions related to events in the summary. DialogSum includes 13k dialogues, with inputs of around 1k tokens and summaries of 130 tokens.
4.2 Online Chat
SAMSum (?) comprises 16k written online dialogues from messaging apps, which linguists craft asked to create conversations similar to those they write daily, reflecting the proportion of topics of their real-life messenger conversations. The messages, each written by one person, contain chit-chats, gossip, arranging meetings, discussing politics, and consulting university assignments. They typically involve two speakers, with an average of 94 tokens per conversation, varying between three and 30 turns  [^7] A turn is the contribution made by a speaker in the form of a single utterance or a statement.. A subset of SAMSum has been adapted into formal third-person language (?) to help models transition from informal dialogue to edited text, addressing the linguistic gap between pre-training and downstream task (detailed in the language challenge, Section 3.1).
FORUM (?) consists of random samples of 100 threads from the online discussion forums on Ubuntu and TripAdvisor, with a total of 556 and 916 posts, respectively. Two human evaluators created summaries of the discussion threads, resulting in two human written summaries per sample.
CRD3 (?) consists of dialogues from the ”Critical Role Dungeon and Dragon” show with summaries collected from the Fandom wiki, featuring 159 episodes with an average dialogue length of 2550 turns.
4.3 Meeting
AMI (?) contains business meetings on the product design process, detailing 137 staged meetings on designing a new remote control. Participants are project managers, marketing experts, user interface, and industrial designers. The dataset includes transcripts and human summaries, with dialogues averaging 6k tokens over 535 turns and four speakers. A modified version, AMI-ITS (?), offers additional annotations for incremental temporal summaries, which provide summaries for 100-second time durations on a subset of AMI.
ICSI (?) has 59 academic group meetings with computer scientists, linguists, and engineers at the International Computer Science Institute (ICSI) in Berkeley, along with their summaries written by hired annotators. The meetings have an average of 819 turns and 13k tokens with six speakers and are about research discussions among students.
QMSum (?) introduces query-based summarization across diverse meeting domains, compiling 1.8k query-summary pairs from 232 meetings, encompassing product design (AMI), academic discussions (ICSI), and committee deliberations. The dialogues feature up to 13k tokens and six speakers. The original task is to summarize the meeting given a stated question. MACSum-Dial (?) is a modification of QMSum, designed for controllable summarization, highlighting mixed attributes such as length, attractiveness, and topic specificity.
ELITR (?) features 120 English technical project meetings in computer science, with each transcript averaging 7k words, 730 turns, and six speakers. The duration of the meetings varies from ten minutes to more than two hours, with an average of one hour long. Meetings shorter than half an hour are exceptions, whereas meetings longer than two hours are topic-oriented mini-workshops, also rather occasional. The original task here slightly differs as a model is required to produce not an abstractive summary but a set of meeting minutes in bullet points.
MeetingBank (?) contains meetings of city councils from six major U.S. cities occurring over the past decade. It contains 1366 meetings spanning 3.6k hours, with a council meeting lasting an average of 2.6 hours and the transcript containing 28k tokens.
4.4 Screenplay
MediaSum (?) contains 463.6k media interview transcripts with brief summaries from National Public Radio (NPR) (?) and CNN, spanning a range of topics. The summaries are based on NPR’s interview overviews and CNN’s topic descriptions, with the latter segmented at commercial breaks to match topics to corresponding interview segments. On average, each transcript in this dataset contains 1.5k words, with summaries of around 11 words, typically involving 7 speakers.
SummScreen (?) consists of 27k instances of TV series transcripts paired with human-written recaps sourced from TV MegaSite and ForeverDreaming. The recaps of ForeverDreaming are based on community contributions stemming from Wikipedia and TVMaze. Transcripts, typically including 28 speakers, average 6.6k tokens, and summaries around 380 tokens.
4.5 Customer Service
TweetSumm (?) contains 1.1k dialogues derived from Twitter customer support exchanges, each with three extractive and three abstractive human-written summaries. Originating from the Kaggle Customer Support On Twitter dataset, these real-world interactions span various industries, such as airlines and retail, averaging 10 turns per dialogue and 36 tokens per abstractive summary.
TODSum (?) is a customer service dataset based on MultiWOZ (?), from which they select the five domains: restaurant, hotel, attraction, taxi, and train. The dataset spans 10k samples with an average dialogue length of 187 utterances and 45 words in the summary.
4.6 Medical
MTS-Dialog (?) is a collection of 1.7k simulated doctor-patient conversations with corresponding clinical notes serving as summaries, sourced from the public Mtsamples collection (?). This dataset encompasses various medical fields, including general medicine, neurology, orthopedics, dermatology, and allergy/immunology, adhering to the SOAP (Subjective, Objective, Assessment, Plan) note format. Dialogues average 18 sentences and 242 words, with summaries typically around 81 words.
4.7 Debates
ADSC (?) features sequences of two-party dialogue chains derived from the Internet Argument Corpus (?), focusing on significant social and political topics like gun control, gay marriage, the death penalty, and abortion. It includes 45 dialogues, each accompanied by 5 unique human-generated summaries, with each summary approximately 150 words in length.
4.8 Others
ConvoSumm (?) consolidates dialogues from four sources: New York Times comments, StackExchange, W3C emails, and Reddit, totaling 2k dialogues with 500 from each domain. Crowdsourced workers from Amazon Mechanical Turk wrote the abstractive summaries with at most 90 tokens. Inputs average 1.1k tokens, with summaries about 70 tokens in length.
FERRANTI (?) is a dataset designed for factual error correction in dialogue summarization, featuring 4k manually annotated items. The original task is to evaluate the factuality of summaries and how to correct these summaries with minimal effort. Drawing on SAMSum and DialogSum, FERRANTI includes summaries produced by models like BART (?) and UniLM (?). Human annotators assess these summaries for accuracy and identify errors, providing a focused tool for improving summary factualness.
DialSum (?) is based on the VisDial dataset (?), where two participants discuss an image from the MSCOCO dataset (?), which features approximately 120k images. Each image thereby has five captions from five different annotators.
4.9 Data Augmentation and Utility
Besides highlighting datasets with a notable community interest, we summarize in this subsection the research on techniques to cope with the data scarcity in dialogue summarization, covering artificial generation techniques of datasets through text generation and curriculum learning strategies to use the available data more effectively.
Artificial Dataset Generation. As discussed earlier, dialogue summarization faces challenges due to the need for adaptability across various domains, structures, and speaker dynamics, which typically would be addressed by training models on diverse datasets (?). However, the scarcity and small size of datasets, with utterances sometimes containing just two to ten turns in areas like customer service and medical, restrict model training for more general applications.
Creating real-world datasets is costly and may conflict with data security and information communication policies, so artificially generated datasets are explored (?). Methods range from simple heuristic-based weak labeling, such as selecting the leading or longest utterance as proxy summary (?), to paraphrasing with updating the summary to maintain coherence (?, ?, ?), to random alterations (e.g., swapping, deletion, dialogue-act-guided insertions) and changing conversation structures (?, ?). The most common method is to use language models to create ground truth summaries either directly (?, ?, ?, ?) or after training them on human summarization techniques through few-shot learning (?).
Techniques to Maximize Dataset Utility. As training Transformer-based models requires extensive data to generalize across various topics and dialogue formats (?), researchers have explored ways to use the limited data available more effectively. A key strategy involves a prompt-based curriculum learning strategy that gradually increases the degree of prompt perturbation (e.g., word swapping, content cutting) to improve the generalization ability of models (?). Another variation is dynamic prompts, which select best-fitting few-shot samples considering dialogue content, size, and speaker number (?). Also explored are prompt transfer techniques from related dialogue domains (?) to bolster the usage of dialogue state information. Prompts are further used to split inputs into domain-invariant and domain-specific content to enhance model generalization (?, ?). Other relevant techniques include freezing most model parameters and training only specific parameters for domain adaption (?, ?, ?).
5 Evaluation
Researchers have adopted metrics from related fields and introduced new ones designed to assess the effectiveness of dialogue summarization methods. These metrics are used as a proxy for quality characteristics, such as coherence, fluency, factuality, and accuracy. In our literature review, we identify 15 metrics used more than twice in 93 papers. We categorize these metrics into four groups: count-based (e.g., ROUGE (?)), model-based (e.g., BARTScore (?)), QA-based (e.g., QAEval (?)), and human evaluation. The automatic metrics strive for alignment with human judgment. However, the correlation between these two is weak for dialogue summarization, and automatic metrics sometimes reward low-quality texts (?). Due to these recognized limitations in fully capturing the nuances of summarization quality, human evaluation is considered the gold standard across most studies. Recent analyses (?, ?) conclude that established automatic metrics may work for a superficial understanding of a model’s performance but do not align well with human judgment in discerning error nuances. These metrics either do not adequately display error occurrence or cannot reflect the impact on the quality of their scores. Hence, a composite metric of count-based, model-based, and QA-based metrics may be required for a more reliable automatic interpretation.
Section 5.1 details the characteristics typically evaluated by the metrics and Sections 5.2, 5.3, 5.4 and 5.5 further summarizes the 15 established metrics in dialogue summarization. Table 1 provides an overview of the identified metrics and their usage throughout the papers considered for this literature review.
5.1 Assessed Characteristics
Metrics assess the quality of a generated text, often in relation to a reference text. Low-quality text may contain more redundancy, incoherence, grammatical errors, poor structure, or inappropriate language, while high-quality text would closely align with the reference. We identify ten key characteristics (in italic) organized into four overarching groups: accuracy, content, readability, and style.
Accuracy is the core of a high-quality summary (?), making adherence to the original text’s truth and, therefore, factuality paramount. A summary must accurately reflect the facts, events, and details from the source without distortion.
Content quality is mainly driven by relevance. A summary must contain the most relevant information, directly addressing the addressee’s informational needs (?). This effort to overcome the salience challenge ensures that summaries prioritize the most important points for the reader. Coverage complements relevance by ensuring that all key topics and arguments from the dialogue are included, offering a comprehensive understanding without significant omissions (?). The goal is to track both the salience and structure challenges with this. Informativeness goes a step further by selecting crucial information that conveys the depth of the input transcript, enabling readers to grasp the main points without needing to refer back to the original text (?).
Readability is essential for making summaries accessible and easy to comprehend. Coherence ensures that information is presented logically with smooth transitions, making the summary easy to follow (?). The structure and organization of a summary further enhances its readability and facilitates information processing, ensuring that it is well-organized and logically constructed (?). Conciseness (?) and non-redundancy (?) are also essential as a summary should contain only the essence of the input transcript without unnecessary details.
Style further contributes to the perceived quality of a text. Consistency in perspective, tense, and style throughout the summary contributes to its readability and professional quality (?). Fluency focuses on grammar, vocabulary, and sentence structure and aims for a natural, easy-to-read summary that is free from phrasing errors, enhancing the information’s overall clarity and accessibility (?).
5.2 Count-Based
Count-based metrics, including N-gram-based measures like BLEU (?), and ROUGE (?), and statistical measures such as Perplexity (?), are often static and rule-based algorithms that have a long history for evaluating text summarization. However, due to their simple approach, they have been criticized for their limitations in capturing overall meaning, fluency, coherence, or factuality in their scores (?).
BLEU (?) measures the precision of generated summaries against reference texts by examining the overlap of word sequences (typically 1- to 3-word N-grams (?)), emphasizing similarity to the original phrasing but risking a linear relationship to noise (?).
ROUGE (?) focuses on lexical similarity (?) and aims to capture the extent to which key content from the source is included in the summary, prioritizing content coverage through recall of N-gram overlap, thus attempting to address BLEU’s limitations by emphasizing content inclusion over mere precision. A frequently used variation is ROUGE-L, which measures the longest common subsequence between a candidate and a reference.
METEOR (?) advances BLEU and ROUGE by incorporating both precision and recall, along with synonym matching for semantic analysis, thus enabling to capture the semantic similarity between a candidate and a reference. Despite its more balanced approach and sentence-level focus, METEOR, like its predecessors, is prone to noise interference (?).
chrF++ (?) further expands the previous metrics by considering precision, recall, and F-score-based N-gram overlap at both character and word levels, offering a more granular analysis (?).
CIDEr (?) is a consensus-based metric that compares the similarity of a generated sentence against a set of human-written reference sentences. The score is an aggregation of cosine similarity scores between the TF-IDF weighted N-grams of the generated and reference sentences, inherently capturing precision, recall, grammaticality, and salience (?, ?, ?).
Perplexity (?) diverges from the previous methods by the statistical approach of gauging a model’s uncertainty in predicting word sequences, with lower scores indicating better alignment with expected language patterns. This measure is, in contrast to the previous, an intrinsic evaluation metric that directly evaluates the language modeling objectives through text predictability rather than assessing the performance on the downstream task. Therefore, additional metrics are required for a holistic evaluation of text generation quality.
5.3 Model-Based
With advancements in language models, there is an increasing focus on model-based evaluation metrics due to their higher correlation with human judgment, though count-based metrics remain popular. Model-based metrics range from those using handcrafted features and learned embeddings to hybrids combining contextual embeddings with count-based approaches. After representing the respective text in a latent space, these metrics consider the similarity between the candidate and reference text, aiming to offer a more nuanced assessment by focusing on semantic similarity, likelihood, and usefulness of summaries. However, despite their potential to adapt to changes in language use (e.g., distribution drifts) (?), these metrics can still be error-prone (?), slower than count-based metrics and may not directly measure the specific characteristics outlined in Section 5.1, making it challenging to discern which particular aspect influences their scores.
BERTScore (?) leverages BERT embeddings (?) to contextually embed the reference and candidate text, calculate the pairwise cosine-similarity on the token level to build a similarity matrix, and greedily pick the highest similarity scores to compute the harmonic mean of precision and recall, thereby capturing semantic nuances (?).
MoverScore (?) was introduced concurrently to BERTScore following a similar approach, as it also leverages BERT embeddings (?) but instead considers the distance between reference and candidate text, hence employing the Word Mover’s Distance (?), a special case of the Earth Mover’s Distance (?), to measure semantic distance. This change in measure allows MoverScore to map semantically related words from one sequence to one word in the other sequence (many-to-one).
BARTScore (?) evaluates the plausibility of generating a reference text from a given generated text, and vice versa, by calculating the log-likelihood of a sequence that a BART (?) model would typically generate based on the given context. This evaluation focuses on assessing both the fluency and the semantic accuracy of the generated content.
BLEURT (?) extends beyond mere embedding comparisons by incorporating a BERT model pre-trained on lexical- and semantic-level supervision signals and fine-tuned on human judgments, enabling it to make detailed assessments of text quality, including coherence and relevance.
BLANC (?) leverages BERT to perform a fill-in-the-blank task, both with and without the generated summary, to assess how informative or helpful the generated text is. The difference in prediction accuracy indicates the utility of the summary in helping understand the text.
FactCC (?) evaluates a summary’s factual consistency with its source document using an entailment classifier, scoring based on the proportion of sentences classified as factual consistent by a BERT model. The metric struggles with complex inferences and subtle nuances beyond direct comparison.
5.4 QA-Based
The QA-based metrics we identify in the literature concentrate on the factual correctness of a summary using external question-answering systems. The effectiveness of these metrics largely depends on the performance of the QA systems, which may not always align with human evaluation standards.
FEQA (?) generates questions based on the summary content and verifies whether their answers can be found in the source document.
SummaQA (?) derives questions from the source text and tries to answer these using the summary. It expands the QA-based evaluation scope by incorporating various question types and emphasizing the summary’s informativeness.
QuestEval (?) combines the approaches of FEQA and SummQA, thereby adopting a bidirectional strategy, generating questions from the summary to compare with the source text and vice versa. This bidirectional evaluation offers a balanced and holistic assessment by considering included and omitted information in the summary.
Type Category Metric Descriptive Tags Usage
Count-based N-gram BLEU word overlap, precision, multiple references 21
ROUGE word overlap, recall, one reference 127
METEOR word overlap, precision and recall, one reference 3
chrF++ character-level, F1 score, one reference 2
CIDEr consensus-based, multiple references 3
[…]
Model-based Hybrid BERTScore token similarity, cosine-similarity, one-to-one 38
MoverScore token similarity, mover distance, one-to-many 3
FactCC entailment classifier, scoring based on consistency 9
[…]
BARTScore mimics human judgement, promptable 13
BLANC fill-in-the-blanc task with and without the summary 2
QA-based FEQA questions based on summary, answered through input 38
SummaQA questions based on input, answered through summary 3
QuestEval combination of FEQA and SummQA 3
Human Evaluation Performance Likert Scale scale, typically 1 (worst) to 5 (best) 6
Pairwise Comparison pick best sample out of two 9
Best-Worst Scaling rank a list of samples 5
[…]
Cohen’s Kappa measure agreements, categorical data, 2 raters 3
Fleiss’ Kappa measure agreement, nominal data, 2 raters 2
Table 1: Relevant metrics and evaluation measures employed in dialogue summarization with more than twice reported use. ’Usage’ states the number of papers reporting the evaluation measure in their publication.
5.5 Human Evaluation
Human evaluation is often considered the gold standard for assessing the quality of a summary. It is typically performed through crowdsourcing annotators (e.g., Amazon Mechanical Turk) where they label samples. We identify established approaches for human evaluations on summary performance and annotator agreement.
Performance. Summary performance is typically evaluated using Likert scales (?, ?, ?, ?), which provide a simple rating system for quality assessment but lack detailed feedback on specific text issues (?). Alternatives like pairwise comparison (?, ?) and best-worst scaling (?, ?) offer more nuanced evaluations, with best-worst scaling noted for its higher reliability (?). Despite these options, the Likert scale remains the predominant method.
Agreement. The reliability of human-generated evaluations hinges on annotator agreement, also referred to as meta-evaluation (?). Despite its significance, the incorporation of agreement metrics is frequently overlooked. We identify three established measures to determine inner-annotator agreements: Krippendorff’s Alpha, Cohen’s Kappa, and Fleiss’ Kappa. Their scores range from 0 (poor reliability) to 1 (perfect reliability), with reported scores typically between 0.65 and 0.85. Krippendorff’s Alpha (?) offers a flexible approach suitable for multiple raters, applicable to any level of measurement, i.e., nominal, ordinal, interval, or ratio, and accommodates both qualitative and quantitative assessments. This measure is particularly valuable with a broad range of variables or when comparing agreements across different measurement scales but assumes that all annotators assess all samples. Cohen’s Kappa (?) is used to measure the agreement between two raters who each classify a set of items into mutually exclusive categories (e.g., yes/no). Fleiss’ Kappa (?) is also tailored for scenarios involving fixed-category classifications but extends Cohen’s Kappa to multiple annotators, excelling in the evaluation of how consistently annotators categorize text segments or dialogue turns into predefined groups.
6 Discussion
6.1 Remarks on Challenges
In this study, we have introduced the CADS taxonomy to organize the inherent challenges of abstractive dialogue summarization (i.e., language, speaker, salience, comprehension, structure, and factuality) and unify their underlying definitions. While presented separately for clarity, the challenges are interdependent and influence each other. Consequently, approaches should be researched to tackle the challenges simultaneously whenever possible. Despite advancements in addressing the challenges, we observe that most of them are still a hurdle for models due to limitations in the Transformer components used in these models, the lack of capabilities in contextualization and few-shot adaption of the encoder-decoder backbone models, and missing mitigation techniques.
Since 2023, NLP has seen a significant shift with the exploration of LLMs and optimized Transformer architecture components (e.g., ring attention (?) and multi-token generation (?)), with their application to dialogue summarization being explored later (?, ?, ?). Initial studies indicate that LLMs match or exceed the performance of task-specific models like DialogLED (?), even with their limitations (e.g., hallucinations, salience) (?). ? (?) show that encoder-decoder models may perform better against the speaker challenge, particularly regarding coreference resolution, and struggle with the structure challenge. LLMs handle the comprehension challenge but seem to be sub-optimal regarding robustness to language and speakers. These early observations state LLMs as a noteworthy alternative to more traditional techniques (e.g., dialogue-style pre-training (?), graph structures to represent speaker structures (?), and special losses tailored to determine salience (?)).
In the following, we discuss how the hurdles of the currently employed encoder-decoder models (identified in Section 3) align with challenges in NLP and discuss techniques introduced so far to mitigate these challenges, thereby deriving how they can aid dialogue summarization.
Bridging the language gap. Adapting established summarization models from structured pre-training data to the less formal nature of dialogue remains challenging. LLMs, especially those trained on non-edited text such as models from the GPT and GEMINI series, are promising in this regard (?). Their few-shot learning capabilities allow for rapid adaptation to new linguistic patterns more effectively. At the same time, the pre-training on a large and diverse corpus improves robustness in handling informal language, style variations, and repetition (?).
Long distance handling and dependency parsing. Recent progress in LLMs such as GPT-4, Claude-3, and Phi-3 shows a significant extension of the processable context length to over 100k tokens, exceeding the context size of earlier models like LED (?), which handle up to 16k tokens. This increase in context allows for handling long dependencies without chunking, which is crucial for understanding complex dialogue structures, speaker dynamics, and content understanding. While traditional models could theoretically handle such lengthy contexts, practical limitations arise from the input-length-dependent quadratic computational costs associated with Transformer attention mechanisms (?). To address these issues, techniques like sparse attention (?), flash attention (training stage) (?) and ring attention (inference stage) (?), sharing weights across attention heads (?), and conditional computation to reduce the overall computational load (?) are proposed to manage large inputs.
Generalizability. Modifying pre-trained models to handle the salience and structure challenges independent of the conversation’s domain is essential for practical applications (e.g., customer service). Established generalization techniques involving prompting (?, ?), few-shot learning (?, ?), and meta-learning (?, ?) are incipient in dialogue summarization. Meanwhile, LLMs build on these techniques to leverage patterns observed during the training stage and boost their generalization capabilities (?). However, LLMs may still struggle with bias mitigation and ensuring broad applicability across various contexts, when ”encountered words are not present in the training corpus or in examples (?, ?, ?).
Coreference resolution. Large-scale pre-trained and fine-tuned language models currently set the benchmark within NLP for coreference resolution (?). Challenges remain with the ambiguities in reference and context (?), and the usage of LLMs for this task.
Hallucination reduction. Avoiding hallucinated content in generated text is a major challenge in NLP tasks using LLMs, especially in culturally sensitive contexts (?, ?). Mitigation techniques include confidence penalty regularization (?, ?) to reduce overconfidence and enhance accuracy.
These advancements throughout NLP suggest upcoming shifts in the key challenges of dialogue summarization.
  • The language challenge can lose impact if few-shot learning is used to prompt LLMs, making it easier to apply document summarization techniques to dialogues directly without additional finetuning.
  • We expect progress for the structure challenge with LLMs as the backbone model, but robustness issues may persist due to issues with generalization and processing long texts.
  • The comprehension challenge, especially in grasping implied meanings, remains largely unsolved and holds significant research potential. Research from related NLP fields such as sentiment analysis (?) suggest that current models still struggle to understand implied content. At the same time, research on retrieval-augmented generation (?) could enhance contextualization, helping models better grasp the context of conversations.
  • For the hurdles implied in the speaker challenge, i.e., coreference resolution and dependency parsing, LLMs are explored but have not yet succeeded due to robustness weaknesses (e.g., errors due to uncommon names, overseeing details in long inputs), signaling that further research is required.
  • The salience challenge currently builds on personalizing summaries and understanding long discussions, which both are gaining attention lately. Due to the novelty, the personalization of summaries requires more research on techniques and benchmarking datasets.
  • The ongoing issue of hallucination in language models continues to pose a challenge in ensuring the factuality in generated summaries.
6.2 Trends in Datasets
Personalized Summarization. In contrast to the typically general summaries, personalized summaries tailored to the users’ needs are becoming increasingly popular (?, ?). Recent advancements have explored cross-attention and decoder self-attention to enhance role-specific information capturing (?). By integrating detailed personal attributes, models can better understand each participant’s background and motivations, leading to more targeted summaries. Dataset-wise, only a few, such as the Chinese CSDS (?), incorporate personalized summaries, whereas established English datasets typically provide only a single, general summary.
Realistic Datasets. As models improve in few-shot learning and do not necessarily require finetuning, the reliance on large-scale datasets is shrinking. Considering this trend, we expect a rise in smaller datasets, which allow for high-quality, realistic examples. These examples may be sufficient to adapt models to new domains, make approaches more robust, and test them against realistic scenarios.
6.3 Trends in Evaluation
Evaluation metrics, vital for indicating performance and comparing new techniques, are currently mostly borrowed from related NLP tasks such as translation or text generation (Section 5). However, the effectiveness of the established metrics (e.g., ROUGE for dialogue summarization as identified in Table 1) is limited (?, ?), and while it can serve as a proxy (?) it provides insufficient insights into the true efficacy of new techniques. The also popular model-based metrics (e.g., BARTScore) seem unable to align well with human judgments (?) for dialogue summarization. Recent developments in NLP use LLM-based metrics such as GEMBA (?) and ICE (?) for evaluation (?), building on LLMs’ advanced text comprehension. This set of metrics thereby mimics human judgment and assesses common aspects (e.g., fluency, coverage, coherence) with continuous (?), Likert scale (?, ?), or probability scores (?). LLM-based metrics offer a promising direction for dialogue summarization evaluation due to their customizability, though this area remains largely unexplored.
7 Final Considerations
7.1 Limitations of Evidence
Given our selection of papers on dialogue summarization, we found limitations in current works that may bias our findings. First, the crawled research studies in dialogue summarization predominantly focus on English-language dialogues, which renders their performance in multilingual or cross-cultural contexts unknown. While we frequently came across Chinese datasets, other languages are underrepresented. As a consequence, we focus our work on English dialogue summarization. Second, most datasets stem from specific sectors such as customer service or business meetings, which may not reflect the variety of dialogue types found in different environments. Lastly, the lack of detailed human evaluation reporting across research works, including inter-annotator agreement scores, further complicates the assessment of new techniques’ usefulness.
7.2 Limitations of Review Process
In our systematic review of dialogue summarization, we encountered methodological challenges that we had to circumvent. First, we focus our search on studies published in English and primarily involving English datasets. While this is a limiting factor, most research on the dialogue summarization task considers only English dialogues, making this a feasible approximation. However, we encourage others to assess the same issues we point out for other languages. Second, despite considering multiple databases, the timing and methods of our search may have missed studies, thus not fully capturing the dialogue summarization field. To mitigate this limitation, we updated our crawl several times throughout writing this article and retrieved works from two databases. We considered both peer-reviewed and non-peer-reviewed papers which risks including less credible studies or misinformation. To mitigate the potential weaknesses of non-peer-reviewed works, we use the ranking from semantic scholar and DBLP, which should account for the paper’s quality and evaluate the studies’ quality based on adherence to established reporting standards of top-tier conferences.
7.3 Conclusion
In this article, we reviewed works on Transformer-based abstractive dialogue summarization. We unified the existing definitions and concepts of dialogue summarization into a taxonomy (CADS) of the field’s six main challenge blocks: language, structure, comprehension, speaker, salience, and reliability-related factuality. We demonstrated how these challenges appear in dialogue summarization and discussed why they occur for established encoder-decoder-based summarization systems. We then grouped and highlighted techniques introduced since 2019 under the challenges they tackle. Despite advancements, we observed that the field is still in its infancy and offers ample opportunities for further research. We also listed datasets used in existing studies to illustrate the extent of data scarcity in dialogue summarization. Additionally, we evaluated common metrics and noted a strong reliance on the ROUGE metric, coupled with a lack of human evaluation reports, which raises doubts about the actual effectiveness of current techniques. In our discussion, we evaluated how current NLP techniques address these challenges and derived implications of using LLMs for the field. We thereby identify a lack of exploring such new models and borrowing effective techniques from related fields to overcome the limitations currently holding back progress. We conclude that while challenges like language may become less relevant, others, such as comprehension and factuality, still require more exploration. We recommend that future work should further discuss and adapt this taxonomy as new challenges, techniques, datasets, and evaluation measures emerge, and as current challenges may be solved naturally through technological innovation. The considered literature can be found in the accompanying repository  [^8] https://anonymous.4open.science/r/LitRev-DialogueSum/ which will be regularly updated.
Acknowledgments
This work was supported by the Lower Saxony Ministry of Science and Culture and the VW Foundation. Frederic Kirstein was supported by the Mercedes-Benz AG Research and Development.
Glossary
LLM
Large Language Model
NLP
Natural Language Processing

### v3

@@
 We cover the main challenges present in dialog summarization (i.e., language, structure, comprehension, speaker, salience, and factuality) and link them to corresponding techniques such as graph-based approaches, additional training tasks, and planning strategies, which typically overly rely on BART-based encoder-decoder models.
-We find that while some challenges, like language, have seen considerable progress, mainly due to training methods, others, such as comprehension, factuality, and salience, remain difficult and hold significant research opportunities.
-We further investigate how these approaches are typically assessed, covering the datasets for the subdomains of dialogue (e.g., meeting, customer service, and medical), the established automatic metrics (e.g., ROUGE), and common human evaluation approaches for assigning scores and evaluating annotator agreement.
-We observe that only a few datasets with community interest span across all subdomains.
-Despite its limitations, the ROUGE metric is the most commonly used, while human evaluation, considered the gold standard, is frequently reported without sufficient detail on inner-annotator agreement and annotation guidelines.
-Additionally, we discuss the possible implications of the recently explored large language models and conclude that despite a potential shift in relevance and difficulty, our described challenge taxonomy remains relevant.
+Recent advances in training methods have led to substantial improvements in language-related challenges.
+However, challenges such as comprehension, factuality, and salience remain difficult and present significant research opportunities.
+We further investigate how these approaches are typically analyzed, covering the datasets for the subdomains of dialogue (e.g., meeting, customer service, and medical), the established automatic metrics (e.g., ROUGE), and common human evaluation approaches for assigning scores and evaluating annotator agreement.
+We observe that only a few datasets (i.e., SAMSum, AMI, DialogSum) are widely used.
+Despite its limitations, the ROUGE metric is the most commonly used, while human evaluation, considered the gold standard, is frequently reported without sufficient detail on the inter-annotator agreement and annotation guidelines.
+Additionally, we discuss the possible implications of the recently explored large language models and conclude that our described challenge taxonomy remains relevant despite a potential shift in relevance and difficulty.
 1 Introduction
-Figure 1: Overview of the big six challenges in dialogue summarization, a short description of the challenges, and the estimation of their progress derived from related sub-challenges.
+Figure 1: Overview of the six challenges in dialogue summarization, including a brief description of each challenge and an estimation of progress for related sub-challenges.
+Progress is evaluated based on two factors: (1) the extent to which mitigation strategies have been developed, and (2) the measurable improvement in summary quality as a result of these strategies.
 Green means mostly mitigated, orange means good progress, and red stands for marked challenges still exist.
 Abstractive dialogue summarization, a task within Natural Language Processing (NLP) and text summarization, entails condensing key information from conversations into succinct and coherent summaries (?).
-This sub-field of text summarization is gaining prominence and is relevant for various real-world scenarios, including customer service (e.g., social media (?) and e-commerce (?)), healthcare (?), daily life (?), meetings (?), and open-domain conversations (e.g., online-chat (?)).
+This sub-field of text summarization is gaining prominence and is relevant for various real-world scenarios, including customer service (e.g., social media, ?, 2021, and e-commerce, ?, 2022), healthcare (?), daily life (?), meetings (?), and open-domain conversations (e.g., online-chat, ?, 2019).
 The relevance of this task, which offers insights into discussed concerns or issues, is thereby driven by the increase in digital conversations in these scenarios (?).
 While the common, manual approach to dialogue summarization is considered to be prone to human errors and consumes significant time and effort, automatic dialogue summarization may effectively reduce this overhead on the user side (?).
-Diverging from traditional text summarization, which typically focuses on formal and linear content such as news articles (?) or scientific publications (?), dialogue summarization confronts unique challenges (?): conversational text is inherently dynamic, interactive, and non-linear (?), marked by verbosity, repetition, and informality.
+Diverging from traditional text summarization, which typically focuses on formal and linear content such as news articles (?) or scientific publications (?), dialogue summarization demonstrates unique challenges (?): conversational text is inherently dynamic, interactive, and non-linear (?), marked by verbosity, repetition, and informality.
 Salient information is often distributed across multiple speakers, mixed with casual and off-topic remarks (?).
 Additionally, the use of elliptical and fragmented sentences, characterized by, e.g., incomplete thoughts and context-specific abbreviations, further hinders the summarization process.
-These complexities make traditional text summarization approaches less effective, which are not trained to bridge the gap between edited text and dialogue and struggle to transfer learned patterns through few-shot learning (?).
+These complexities make traditional text summarization approaches less effective, as they are not trained to bridge the gap between edited text and dialogue and struggle to transfer learned patterns through few-shot learning (?).
 Despite advancements achieved by leveraging and adapting general-purpose pre-trained neural language models (?, ?), and the recent trend to explore Large Language Models (LLMs) for this task (?), there remains a notable gap in the field’s foundational understanding of the challenges when processing conversation transcripts.
 While research papers introduce techniques addressing similar challenges, their definitions and interpretations of these challenges can vary markedly.
-For instance, while some describe the informal and ungrammatical nature of spoken language as a language challenge (?, ?), others frame the challenge as different individual styles and unstructured expressions (?, ?).
-Both points of view are thereby valid but, on their own, focus just on a subpart of the language challenge and blur the understanding of it.
+For instance, while some describe the informal and ungrammatical nature of spoken language as a language challenge (?, ?), others frame the language challenge as different individual styles and unstructured expressions (?, ?).
+Both viewpoints are thereby valid but, on their own, focus just on a subpart of the language challenge and blur the complete understanding of it.
 Current surveys leave out the challenges in dialogue summarization and their variations in definitions, focusing instead on details on datasets, metrics, and techniques.
-As a consequence, there is no clear understanding of which aspects of dialogue summarization are well-addressed and where significant research potential lies.
-Our systematic review consolidates the existing knowledge researched in the field considering Transformer-based models, aiming to provide a comprehensive overview of the challenges and their progress.
-We further consider the center interest in LLMs and discuss their impact on the relevance of the dialogue summarization challenges.
+Consequently, there is no clear understanding of which aspects of dialogue summarization are well-addressed and where significant research potential lies.
+Our systematic review consolidates the knowledge researched in the field considering Transformer-based models, aiming to provide a comprehensive overview of the challenges and their progress.
+We further consider the interest in LLMs and discuss their impact on the relevance of the dialogue summarization challenges.
 This article focuses on four key areas, namely challenges, dataset overviews, techniques, and evaluation methods.
-In challenges, we categorize known challenges for Transformer-based models found when dealing with dialogue transcripts into six broader challenge blocks, i.e., language, structure, comprehension, speaker, salience, and factuality.
+In challenges, we categorize known hurdles for Transformer-based models found when dealing with dialogue transcripts into six broader challenge blocks, i.e., language, structure, comprehension, speaker, salience, and factuality.
 Our proposed taxonomy is displayed in Figure 1, including the six challenge blocks, a short description, and exemplary sub-challenges of the challenge blocks.
@@
 In techniques of dialogue summarization, we then match 97 techniques proposed since 2019 to the six challenge blocks.
-Furthermore, in evaluation methods, we give an overview of typical evaluation metrics used, spanning count-based metrics (e.g., ROUGE (?)), model-based (e.g., BERTScore (?)), QA-based (e.g., QuestEval (?)), and human evaluation metrics for performance and annotator agreement.
+Furthermore, in evaluation methods, we provide an overview of typical evaluation metrics used, spanning count-based metrics (e.g., ROUGE, ?, 2004), model-based (e.g., BERTScore, ?, 2020), QA-based (e.g., QuestEval, ?, 2021), and human evaluation metrics for performance and annotator agreement.
 This systematic review contributes to a better understanding of the dialogue summarization problem, unifies the inherent definitions, overviews established techniques, and demonstrates the scarcity of datasets and fitting evaluation metrics.
@@
 In Section 3, we provide the general problem definition and our challenges taxonomy consisting of six challenge blocks inherent in conversation scripts, further linking proposed approaches to handle the challenges.
-Section 4 overviews prominent datasets while the display of evaluation approaches in Section 5 helps readers choose suitable indices to evaluate the effectiveness of a model.
+Section 4 overviews prominent datasets, while the display of evaluation approaches in Section 5 helps readers choose suitable indices to evaluate the effectiveness of a model.
 Finally, Section 6 discusses future research directions followed by conclusions in Section 7.
 All resources for our review are publicly available.
-[^1] https://anonymous.4open.science/r/LitRev-DialogueSum/
+[^1] https://github.com/FKIRSTE/LitRev-DialogueSum
 1.1 Existing Reviews in Dialogue Summarization
@@
 ?
-(?) analyze dialogue summarization techniques up to early 2022, categorizing techniques by conversational context and component modeling (e.g., speaker, addressee, and utterance modeling to solve the ’Who says what to whom’ paradigm), covering a subset of the conversation challenges indirectly.
+(?) analyze dialogue summarization techniques up to early 2022, categorizing techniques by conversational context and component modeling (e.g., speaker, addressee, and utterance modeling to solve the ‘Who says what to whom’ paradigm), covering a subset of the conversation challenges indirectly.
 ?
@@
 We analyze datasets (Section 4) and evaluation metrics (Section 5), linking them to the identified challenges.
-Additionally, we discuss the implications of the shift from smaller encoder-decoder models such as BART (?) to LLMs and the implications for the relevance of the challenges with this new backbone model type (Section 6).
-Our review concentrates exclusively on dialogue summarization and does not encompass broader areas such as general text summarization (?) or dialogue generation (?).
+Additionally, we discuss the shift from smaller encoder-decoder models such as BART (?) to LLMs and the implications for the relevance of the challenges with this new backbone model type (Section 6).
+Our review focuses exclusively on dialogue summarization and does not encompass broader areas such as general text summarization (?) or dialogue generation (?).
 2 Methodology
 We use the PRISMA checklist (?) for organizing our literature review, as it is an established approach to set up a comprehensive systematic literature review while reducing the potential for incomplete data and biases in content selection and presentation (?).
-This ensures that our approach is structured, transparent, and reproducible.
-Figure 2: An overview of the different stages of our methodology and their inherent sub-stages.
-Our methodology comprises two stages as detailed in Figure 2.
+This checklist ensures that our approach is structured, transparent, and reproducible.
+Figure 2: An overview of the different stages of our methodology and their inherent sub-stages. The red boxes indicate the number of papers considered in the respective step.
+Our methodology comprises two stages, as detailed in Figure 2.
 Retrieval Stage.
@@
 We evaluate 1262 unique papers from 2019 until 2024, offering an up-to-date perspective on the state of the art in dialogue summarization using Transformer-based architectures.
-We decide against including references from selected papers that did not appear in our crawl as additional sources to maintain an unbiased selection process and avoid complicating the stopping criterion for literature inclusion.
+We decided against including references from selected papers that did not appear in our crawl as additional sources to maintain an unbiased selection process and avoid complicating the stopping criterion for literature inclusion.
 Considering our exclusion criteria, such as multilingualism, non-abstractive, and non-Transformer-based approaches, we select 133 papers for our literature review.
@@
 While we acknowledge the potential for these databases to miss the most recent uploads, our strategy minimizes this risk.
-We exclude databases such as Google Scholar, which overly returned grey literature, and Web of Science, which emphasized extractive summarization techniques during our pre-testing, to maintain a focus on peer-reviewed literature in abstractive dialogue summarization.
+We exclude databases such as Google Scholar, which overly returned gray literature, and Web of Science, which emphasized extractive summarization techniques during our pre-testing, to maintain a focus on peer-reviewed literature in abstractive dialogue summarization.
 We design an automatic pipeline to retrieve query-related works and remove duplicates.
@@
 Our search strategy uses key terms related to dialogue summarization identified during our initial related work analysis (Section 1.1).
-Anchor terms such as ’dialogue summarization’ and ’multi-party conversation’ are combined with specific search terms such as ’technique’.
-We use four anchor terms and 28 search terms, resulting in 112 search queries for each database.
-We present the complete list and search keywords in our repository [^5] https://anonymous.4open.science/r/LitRev-DialogueSum/ .
+Anchor terms such as ‘dialogue summarization’ and ‘multi-party conversation’ are combined with specific search terms such as ‘technique.’ We use four anchor terms and 28 search terms, resulting in 112 search queries for each database.
+We present the complete list and search keywords in our repository [^5] https://github.com/FKIRSTE/LitRev-DialogueSum .
 Employing a Python-based pipeline (Section 9), we systematically extract documents from Semantic Scholar and DBLP APIs, merging and unifying the results into tabular data.
 Between January 2019 and March 2024, we retrieved 732 publications from DBLP and 968 from Semantic Scholar.
-After removing 438 duplicates identified across both sources, we have 1262 unique publications at the end of the stage.
+After removing 438 duplicates identified across both sources, we conclude the stage with 1262 unique publications.
 These results are tagged with their respective queries and compiled into a CSV file for subsequent selection and analysis.
@@
 For our manual filtering process, we define eligibility criteria to ensure the focus of our literature review.
-We look at long and short, open-access papers published since 2019 that align with the rise of Transformer-based text summarization models such as BART (?) and PEGASUS (?), which have been the core model for most approaches in the field.
-This timeframe further allows us to concentrate on current challenges rather than including some that may have lost relevance (e.g., negation (?)) because of new advancements in the field.
-We exclude non-English primary datasets, multi-modal studies, and extractive or non-Transformer-based methods to maintain a focus on English-language research on abstractive dialogue summarization.
+We consider open-access papers published since 2019 that align with the rise of Transformer-based text summarization models such as BART (?) and PEGASUS (?), which have been the core model for most approaches in the field.
+This timeframe further allows us to concentrate on current challenges rather than including some that may have lost relevance (e.g., negation, ?, 2021) because of advancements in the field.
+We exclude non-English primary datasets, multi-modal studies, and extractive or non-Transformer-based methods and focus on English-language research on abstractive dialogue summarization.
 Manual Paper Processing.
@@
 They then review the full texts, checking if the works still meet the eligibility criteria, and decide whether to include, exclude, or discuss each paper.
-Disagreements are resolved through consensus.
-They also track features such as used datasets, employed backbone models, proposed techniques, and evaluation approaches.
+Each paper undergoes a secondary review to minimize bias, reaffirming initial selections and categorizations, thereby enhancing the study’s integrity.
+Eventual disagreements are resolved through consensus.
+The annotators also track features such as used datasets, employed backbone models, proposed techniques, and evaluation approaches.
 This step filters out 53 papers, leaving 133 to compose the literature review and detail challenges, techniques, datasets, and evaluation metrics.
-To further minimize bias, a secondary review is performed as needed, reaffirming initial selections and categorizations, thereby enhancing the study’s integrity.
 3 Challenges and Progress
 Abstractive dialogue summarization aims to condense multi-party conversations into their key points, ranging from casual chats to expert discussions, and whether these points were explicitly stated or inferred (?).
-The task was first thoroughly defined in 2002 by ?
-(?), outlining challenges unique to spoken dialogues such as handling speech disfluency, the lack of clear sentence boundaries, distribution of salient information across various speakers and turns, resolving references, identifying discourse structures, and dealing with inaccuracies caused by speech recognition errors.
-While some definitions take a mathematical approach (?), viewing dialogues as sequences of turns that are compressed and each associated with a specific speaker, others adopt a broader description of the task as ”extracting useful information from a dialogue” (?).
-Although all definitions contain relevant, overlapping aspects, they lay down different viewpoints, each considering a different set of challenges to define dialogue summarization.
+The task was first thoroughly defined in 2002 by ?, outlining challenges unique to spoken dialogues such as handling speech disfluency, the lack of clear sentence boundaries, distribution of salient information across various speakers and turns, resolving references, identifying discourse structures, and dealing with inaccuracies caused by speech recognition errors.
+While some definitions take a mathematical approach (?), viewing dialogues as sequences of turns that are compressed and associated with a specific speaker, other definitions adopt a broader description of the task as ‘extracting useful information from a dialogue’ (?).
+Although all definitions contain relevant, overlapping aspects, they have different viewpoints, each considering a different set of challenges to define dialogue summarization.
 Our goal is to organize these definitions and provide a concise, structured overview of the progress in abstractive dialogue summarization, extending the earlier challenge-based concept (?) and proposing the Challenges of Abstractive Dialogue Summarization (CADS) Taxonomy (Figure 1) tailored to the current Transformer-based model architecture.
-This taxonomy consists of six challenges: language, structure, comprehension, speaker, salience, and factuality.
-The first five challenges are thereby related to the input, while the last one concerns reliability.
+This taxonomy comprises six challenges: language, structure, comprehension, speaker, salience, and factuality.
+The first five challenges are related to the input, while the last concerns reliability.
+To develop this comprehensive framework, we employed human annotators (Section 2.2) to systematically analyze and categorize recurring challenges in the literature and consolidate them into the six challenges.
+Each pillar encompasses related sub-challenges that represent key focus areas consistently addressed in the literature along the broader challenges.
 In the following Sections 3.1, 3.2, 3.3, 3.4, 3.5 and 3.6, we detail the CADS Taxonomy as shown in Figure 1.
-For this section, we consider 91 papers in total, from which 73 papers define the challenge characteristics (sub-challenges in italic) and 74 describe explored techniques.
+An overview of the approach strategies proposed to mitigate the individual challenges is shown in Table 1.
+Table 2 summarizes the expected errors when challenges are not successfully mitigated, using the analysis by ?
+(?) as a base for the individual error types.
+For this section, we consider 91 papers, from which 73 papers define the challenge characteristics (sub-challenges in italic) and 74 describe explored techniques.
 We exclude papers mainly discussing evaluation strategies and datasets from this subset and cover them in Sections 4 and 5.
+Challenge Approach Category Related Papers
+Language Pre-training ? (?, ?, ?, ?)
+Training tasks ? (?, ?, ?, ?, ?)
+Pre-processing ? (?)
+Structure Pre-training ? (?, ?, ?)
+Training tasks ? (?, ?, ?, ?)
+Architecture modification ? (?, ?, ?, ?, ?)
+Importance measures ? (?, ?)
+Comprehension Architecture modification ? (?)
+Speaker Training tasks ? (?, ?, ?, ?)
+Architecture modification ? (?, ?, ?, ?)
+Pre-processing ? (?, ?)
+Post-processing ? (?, ?)
+Salience Pre-training ? (?, ?)
+Training tasks ? (?, ?, ?)
+Architecture modification ? (?, ?)
+Human feedback ? (?)
+Loss function ? (?)
+Pre-processing ? (?, ?)
+Factuality Training tasks ? (?, ?)
+Architecture modification ? (?, ?, ?, ?)
+Human feedback ? (?)
+Loss function ? (?, ?)
+Post-processing ? (?, ?)
+Table 1: Overview of challenges, major approach trends, and corresponding literature. More details are stated in the article’s accompanying repository:
+https://github.com/FKIRSTE/LitRev-DialogueSum.
 3.1 Language
@@
 The language challenge describes the idiosyncratic nature of spoken language and individual speech patterns.
-This includes linguistic subtleties such as informal expressions (e.g., ’yeah,’), ungrammatical structures, colloquialisms, personal vocabulary, and linguistic noise (e.g., filler words) (?, ?, ?, ?, ?).
-It further covers content repetition (i.e., speakers restate or rephrase information for emphasis or clarity) (?, ?, ?), and domain-specific terminology (e.g., medical terms) (?, ?, ?, ?).
+It includes linguistic subtleties such as informal expressions (e.g., ‘yeah,’), ungrammatical structures, colloquialisms, personal vocabulary, and linguistic noise such as filler words (?, ?, ?, ?, ?).
+It further covers content repetition, i.e., speakers restate or rephrase information for emphasis or clarity (?, ?, ?) and domain-specific terminology, e.g., medical terms (?, ?, ?, ?).
 These elements demand that models accurately interpret and adapt to the linguistic features of dialogues (?).
@@
 Failing to handle this challenge may result in a loss of coherence and clarity, repetition of content, and factuality issues.
-Figure 3: Example dialogue to demonstrate examples of the idiosyncratic nature of spoken language and individual speech patterns.
-Red displays disfluencies, blue highlights personal speech patterns, orange displays the colloquialism, green represent informal expressions
+Figure 3: Dialogue snippet showing examples of the idiosyncratic nature of spoken language and individual speech patterns.
+Red displays disfluencies, blue highlights personal speech patterns, orange stands for colloquialism, green represent informal expressions.
 Approaches.
-Transformer-based models often struggle with the language characteristics due to a gap between their pre-training data, typically well-edited texts (e.g., news articles, research papers, Wikipedia entries) (?), and the characteristics of spoken dialogue not reflected in these edited texts (?).
+Transformer-based models often struggle with the language characteristics due to a gap between their pre-training data, typically well-edited texts (e.g., news articles, research papers, Wikipedia entries, ?, 2020), and the characteristics of spoken dialogue not reflected in these edited texts (?).
 The transfer between these text styles is hindered by the scarcity of diverse, dialogue-oriented pre-training data, resulting in a lack of exposure to dialogue data during training, consequently reducing models’ performance.
@@
 Current research aims to bridge this gap between the formal language of the pre-training data and dialogues by retraining models with dialogue-focused tasks, considering both single- and multi-task setups.
-Explored tasks span masking key dialogue elements (e.g., pronouns entities, high-content tokens, and words) (?) and part-of-speech tagging (?).
+Explored tasks span masking key dialogue elements (e.g., pronouns entities, high-content tokens, and words, ?, 2021) and part-of-speech tagging (?).
 More recent tasks aim to transform informal, first-person dialogue into a structured, third-person narrative through changing speaker names, adjusting grammar, and adding emotional context (?), or simulating a dialogue by modifying pre-training documents into a conversation structure (?).
-Teaching models to understand the ’who-did-what’ structure through a specific task is further gaining traction (?).
-Other explored strategies include pre-processing the input for anaphora resolution (?) and adapted pre-training (?, ?, ?).
+Teaching models to understand the ‘who-did-what’ structure through a specific task is further gaining traction (?).
+Other explored strategies include pre-processing the input for anaphora resolution (?) and adapted pre-training, e.g., to adjust to specific terminology (?, ?, ?).
 3.2 Structure
 Characteristics.
-Aiming at extracting the built-up of a conversation, the structure challenge revolves around segmenting a dialogue transcript and creating an ordered summary based on the extracted structures (?).
-This challenge involves identifying the topic flow within a conversation (?, ?, ?, ?), tracking the dialogue phases (e.g., problem identification, decision making) (?, ?)), and analyzing the utterance interaction (i.e., relationship and dependencies between utterances) (?, ?, ?).
-In Figure 4, we show a dialogue snippet exhibiting topic flow, dialogue phases, and utterance interactions.
-This can result in summaries lacking coherence, completeness, and depth (?, ?, ?, ?, ?).
+Aiming to extract a conversation’s built-up, the structure challenge is about segmenting a dialogue transcript and creating an ordered summary based on the extracted structures (?).
+This challenge involves identifying the topic flow of a conversation (?, ?, ?, ?), tracking the dialogue phases (e.g., problem identification, decision making, ?, 2021; ?, 2023), and analyzing the utterance dependency (i.e., relationship and dependencies between utterances, ?, 2021a; ?
+2021b; ?, 2022).
+The dialogue in Figure 4 shows topic flow, dialogue phases, and utterance dependency.
+This challenge can lead to summaries lacking coherence, completeness, and depth (?, ?, ?, ?, ?).
 Figure 4: Excerpt from a conversation to demonstrate the structural challenge, with a detailed view on one topic out of multiple topics touched in the whole conversation.
@@
 Approaches.
-Transformer-based models face two primary hurdles when handling the structural challenge, i.e., adapting to the variability in dialogue structures, such as differing topic sequences and conversation phases depending on the conversation type, and managing long-range dependencies to link utterances.
-The first difficulty is related to the lack of varied datasets across different dialogue types on which models can train to improve generalization as well as that most pre-trained models used are expecting clearer structural signals due to the more structured data used during pre-training (Section 3.1).
-The second problem relates to a weakness of transformed-based architectures: their limitation in handling long sequences, which stems from the practical limitation of a fix-sized context window (?), the diluted attention on a single token when many tokens are processed (?), and the bias towards recent information (?).
+Transformer-based models face two primary hurdles when handling the structural challenge.
+First, they must adapt to the variability in dialogue structures, such as differing topic sequences and conversation phases depending on the conversation type.
+Second, they need to track and connect long-range dependencies to link utterances.
+The first difficulty is related to the lack of varied datasets across different dialogue types on which models can train to improve generalization as well as that most pre-trained models used are expecting more precise structural signals due to the more structured data used during pre-training (Section 3.1).
+The second problem relates to a weakness of transformed-based architectures: their limitation in handling long sequences, which stems from the practical limitation of a fix-sized context window (?), the blurred attention on a single token when many tokens are processed (?), and the bias towards recent information (?).
 Recent methods employ graph structures to understand and use dialogue structures more effectively.
@@
 Abstract Meaning Representation (AMR) graphs are employed to capture overarching themes (?), detailed sentence-level connections, and entity interactions, enhancing content comprehension (?).
-Meanwhile, a hybrid approach combining traditional language models with graph neural networks captures the utterance interactions and addresses the long-range dependency sub-challenge (?).
+Meanwhile, a hybrid approach combining traditional language models with graph neural networks captures the utterance dependency and addresses the long-range dependency sub-challenge (?).
 Similarly, ConceptNet (?) is used for thematic linking (?).
 To manage long dialogue transcripts beyond context-size limits and the inherent long-distance relations, researchers are moving away from simply truncating inputs to context length to avoid losing vital information.
-A leading method segments the summarization process into stages (e.g., topics and dialogue phases) for coherent processing (?, ?, ?).
+An established approach segments the summarization process into stages (e.g., topics and dialogue phases) for coherent processing (?, ?, ?).
 The segmentation and multi-pass strategy iteratively summarizes parts of the dialogue and combines partial summaries with a subset of the remaining content in each iteration to build a complete summary (?).
-The $Summ^{N}$ framework (?), summarizes dialogue segments independently before refining them together, offering a detailed and scalable solution.
-Other techniques involve additional training tasks (?, ?, ?, ?), sentence importance measures to group sentences into sub-topics (?, ?), while hierarchical setups to differentiate dialogue parts based on who’s speaking (word level) and what’s being discussed (turn level) are also popular (?).
+The $Summ^{N}$ framework (?) summarizes dialogue segments independently before refining them together, offering a scalable solution.
+Other techniques involve additional training tasks (?, ?, ?, ?), sentence importance measures to group sentences into sub-topics (?, ?), while hierarchical setups to differentiate dialogue parts based on who is speaking (word level) and what is being discussed (turn level) are also popular (?).
+Figure 5: Example conversation to demonstrate the comprehension challenge including direct and implied content.
+Directly stated (orange) are the explicit updates by Bob and Charlie about their respective tasks.
+Green marks implied content such as a deadline for a demo and a current focus on performance and finalizing features (unmentioned background knowledge), Redis implies an understanding of its role in performance enhancement (organizational knowledge and reference to prior discussions on data requirements).
 3.3 Comprehension
 Characteristics.
-The comprehension challenge involves accurately understanding and contextualizing utterances so models can grasp and leverage these insights for summarization (?).
+The comprehension challenge involves accurately understanding and contextualizing utterances such that models can grasp and leverage these insights for summarization (?).
 This includes direct and implied content, e.g., unmentioned background knowledge, organizational knowledge, and prior discussions not explicitly referred to (examples in Figure 5).
 To infer and understand this content, the whole context of a dialogue is required, spanning both the local context, focusing on neighboring words and sentences, and the global context, covering the entire dialogue’s informational flow (?).
-Figure 5: Example conversation to demonstrate the comprehension challenge including direct and implied content.
-Directly stated (orange background) are the explicit updates by Bob and Charlie about their respective tasks.
-Green marks implied content such as a deadline for a demo and a current focus on performance and finalizing features (unmentioned background knowledge), Redis implies an understanding of its role in performance enhancement (organizational knowledge and reference to prior discussions on data requirements).
+Insufficient comprehension can lead to misunderstanding the content, omitting parts, or falsely presenting the information.
 Approaches.
-Transformer-based models struggle with comprehension and contextualization due to the underlying dependency parsing challenges, particularly with unclear sentence boundaries as they appear in spoken language, which can lead to catastrophic forgetting and result in instability in handling long-range dependencies (?).
-Additionally, these models often interpret the text literally, affecting their ability to grasp nuanced meanings (?, ?).
+Transformer-based models struggle with comprehension and contextualization due to the underlying dependency parsing challenges, particularly with unclear sentence boundaries as they appear in spoken language, which can lead to catastrophic forgetting and result in unreliability in handling long-range dependencies (?).
+Additionally, these models interpret the text literally, affecting their ability to grasp nuanced meanings (?, ?).
 As these models often rely on syntactic relationships between words, they miss the nuanced understanding humans use to interpret implied meanings of sarcasm, irony, euphemisms, and similar pragmatic elements (?).
@@
 Participants, their interactions, and how they influence the dynamics of a conversation are part of the speaker challenge.
-This includes the individual participant roles (e.g., ”project manager”, ”customer service agent”) (?, ?), topic-dependent role changes (?, ?, ?, ?, ?, ?, ?), their interactions (e.g., contributions depending on viewpoints (?, ?, ?), backgrounds (?, ?, ?)), and references to entities and other participants (?, ?, ?, ?).
+This includes the individual participant roles (e.g., ‘project manager,’ ‘customer service agent’, ?, 2021b; ?, 2022), topic-dependent role changes (?, 2021; ?, 2022a; ?, 2022; ?, 2022; ?, 2022; ?, 2022; ?, 2023), their interactions (e.g., contributions depending on viewpoints, ?, 2021; ?, 2022; ?, 2022; and on backgrounds, ?, 2021; ?, 2021; ?, 2022), and references to entities and other participants (?, ?, ?, ?).
 Examples of the speaker characteristics are shown in Figure 6.
 Failing to capture these speaker characteristics may lead to missing context, misaligned entity associations, and, ultimately, incorrect factual summaries.
-The significance of examining speaker dynamics and roles is well-recognized in other fields such as linguistics (?) and is, together with the language challenge of the preceded analysis, a common hook to demonstrate and motivate techniques for dialogue summarization.
+The significance of examining speaker dynamics and roles is well-recognized in other fields such as linguistics (?) and is, together with the language challenge of the preceded analysis (Section 3.1), a typical hook to demonstrate and motivate techniques for dialogue summarization.
 Figure 6: Example conversation of the speaker challenge with defined roles (Alex as customer, Jordan as customer support agent, Taylor as technical specialist).
 The dialogue shows a change of roles when Jordan calls Taylor (marked with orange), shifting from Jordan as a primary troubleshooter to a facilitator for Taylor’s technical interventions.
-Further, multiple direct references to entities highlighted with blue (e.g., ”device”, ”error codes”) and indirect references to participants in red (e.g., ”you”) are incorporated.
+Further, multiple direct references to entities highlighted with blue (e.g., ‘device’, ‘error codes’) and indirect references to participants in red (e.g., ‘you’) are incorporated.
 Approaches.
-The speaker challenge arises from identifying participants, tracking their actions and entity references, and revealing participant relationships.
-These challenges can be aligned with the difficulties of named entity recognition (NER), coreference resolution, and dependency parsing.
-Although advancements with Transformer-based models regarding NER, the explored techniques struggle to accurately identify and categorize entities across different text types and domains and with unstructured texts (?, ?, ?), rendering realistic dialogue scenarios as a challenging field.
-Coreference resolution poses difficulties for techniques in accurately identifying mentions and contextual links (?) and understanding the linguistic structures and relationships (?), extending the understanding challenge (Section 3.3).
+The speaker challenge stems from identifying participants, tracking their actions and entity references, and revealing participant relationships.
+These aspects correspond to the difficulties of named entity recognition (NER), coreference resolution, and dependency parsing.
+Despite advancements with Transformer-based models regarding NER, current techniques struggle with entity identification and categorization, particularly in unstructured texts and across diverse domains (?, ?, ?).
+This limitation is especially pronounced in realistic dialogue scenarios.
+Coreference resolution poses difficulties for techniques in accurately identifying mentions and contextual links (?) and understanding the linguistic structures (?), extending the understanding challenge (Section 3.3).
 Works typically use dense vectors (?, ?) to accurately represent speaker roles in dialogue and act as labels for utterances (?, ?), offering insights into each speaker’s function within the dialogue.
-Speaker vectors are either randomly initialized and trainable (?, ?) or produced by small neural networks (?, ?) and represent different roles, e.g., ”industrial designer” (?), ”judge” (?), which are appended to the embedding of the speaker’s turn.
-For capturing speaker dynamics and resolving coreferences, graph-based models are an established approach, e.g., to build a graph representing each speaker’s main ideas and their discourse, alongside their interactions, reflecting inner- and inter-speaker structures (?, ?).
+Speaker vectors are either randomly initialized and trainable (?, ?) or produced by small neural networks (?, ?) and represent different roles, e.g., ‘industrial designer’ (?), ‘judge’ (?), which are appended to the embedding of the speaker’s turn.
+For capturing speaker dynamics and resolving coreferences, graph-based models are an established approach, e.g., representing each speaker’s main ideas and their discourse, alongside their interactions, reflecting inner- and inter-speaker structures (?, ?).
 Some other popular techniques involve enhancing Transformers’ self-attention to analyze intra- and inter-speaker dynamics (?, ?), additional training tasks (?, ?, ?), and pre-and post-processing techniques to replace pronouns and correct mistakes (?, ?, ?, ?) for coreference resolution.
+Figure 7: Example of the salience challenge with a conversation being summarized for a general audience and tailored summaries for participants with a specific focus on content relevant to them.
+Figure 8: Example of the factuality challenge where the generated summary exhibits extrinsic errors through hallucinating events (blue background) such as the finalization of a marketing strategy and a detailed review of financial strategies.
+The intrinsic error reporting the wrong date of the meeting is marked with red.
 3.5 Salience
@@
 Figure 7 demonstrates the differences between personalized and generally applicable summaries, as the generated texts for Simon and Nadia are tailored to their role, stands, and tasks, whereas the general summary provides a neutral overview.
-Failure to effectively locate the scattered information or understand what is considered salient can impact the depth and relevance of the generated summaries, focusing on unimportant content while omitting relevant ones (?, ?, ?, ?).
-Figure 7: Example of the salience challenge with a conversation being summarized for a general audience and tailored summaries for participants with a specific focus on content relevant to them.
+In contrast to traditional summarization, where salience may be related to the position of sentences (positional salience), dialogues are more dynamic, and salience emerges from how ideas evolve over the conversation.
+Failure to effectively locate the scattered information or understand what is considered salient can impact the depth and relevance of the generated summaries, focusing on unimportant content while omitting relevant information bits (?, ?, ?, ?).
 Approaches.
@@
 To better determine salient content for specific participants, works propose a fine-tuning task in which the topic or user perspective to be focused on is passed as an input along the dialogue text (?), apply question-driven pre-training (?) and use dynamic prompts to direct model attention to key dialogue aspects (?).
-Additional strategies involve direct text manipulation (?), personal named entity planning to concentrate on specific entities (?), adapting attention mechanisms (?), and integrating human feedback (?) or graph structures (?) to further tailor summaries to the dialogue’s essential parts.
+Additional strategies involve direct text manipulation (?), personal named entity planning to concentrate on specific entities (?), adapting attention mechanisms (?), and integrating human feedback (?) or graph structures (?) to further tailor summaries to the dialogue’s core information.
+Challenge Related Errors
+Language incoherence, repetition, hallucination
+Structure incoherence, omission, lack of depth
+Comprehension lack of context understanding, omission, hallucination
+Speaker lack of context understanding, false coreference resolution, hallucination
+Salience irrelevance, lack of context understanding
+Factuality hallucination
+Table 2: Definition of eight to-be-expected error types in dialogue summarization, based on existing meeting summarization related error types (?).
 3.6 Factuality
 In contrast to the previous input-related challenges, the factuality challenge describes the problem of correcting false content in a generated summary caused by a non-robust architecture.
-Factual incorrect content contains intrinsic errors, such as hallucinating events not present in the original transcript, and extrinsic errors, such as incorrect coreference resolution misrepresenting actual event details, and wrong conclusions stemming from negation (?).
-Intrinsic and extrinsic errors are shown in a sample dialogue in Figure 8.
-Given the importance of ensuring dialogue summarizers accurately reflect the conversation, research focuses on creating safety measures to maintain summary credibility despite these challenges.
-Figure 8: Example of the factuality challenge where the generated summary exhibits intrinsic errors through hallucinating events (blue background) such as the finalization of a marketing strategy and a detailed review of financial strategies.
-The extrinsic error reporting the wrong date of the meeting is marked with red.
+Factual incorrect content contains extrinsic errors, such as hallucinating events not present in the original transcript, and intrinsic errors, such as incorrect coreference resolution misrepresenting actual event details, and wrong conclusions stemming from negation (?).
+Extrinsic and intrinsic errors are shown in a sample dialogue in Figure 8.
+Given the importance of ensuring summaries accurately reflect the conversation, research focuses on creating safety measures to maintain credibility.
 Approaches.
 Research interest predominantly targets post-processing approaches that rewrite the generated summary to correct factual errors.
-Notable approaches here are the usage of LLMs for error detection and correcting (?) and reformulating the summary to closely match the original dialogue’s predictive capabilities for subsequent content (?).
-Other approaches explore the use of auxiliary tasks (?) to estimate the factual aspects and predict the missing aspects in a summary (?), losses such as encouraging a model to generate sentences about content not yet covered and differentiating factual from non-factual sentences (?, ?), and human feedback (?).
-Further, architecture-modifying approaches are explored, such as a new encoder to grasp dialogue states (?) or to handle graph structures, where events of dialogue are captured and organized in a graph.
-A special attention mechanism is used to fill in summary (?), and a hierarchical approach builds on sub-summaries (?).
-Another noteworthy technique is the generation of a sketch to structure out the final summary along this plan (?).
-Negation is a long-standing challenge for language models often neglected in computational studies (?, ?).
+Notable approaches are the usage of LLMs for error detection and correcting (?) and reformulating the summary to closely match the original dialogue’s predictive capabilities for subsequent content (?).
+Other approaches explore the use of auxiliary tasks (?) to estimate the factual aspects and predict the missing aspects in summary (?), losses such as encouraging a model to generate sentences about content not yet covered and differentiating factual from non-factual sentences (?, ?), and human feedback (?).
+Further, architecture-modifying approaches are explored, such as a new encoder to grasp dialogue states (?) or to handle graph structures, where dialogue events are captured and organized in a graph.
+A slot-driven beam search algorithm is used in a filling-in-the-blanks setup to give priority to generating salient elements in the summary (?), and a hierarchical approach builds on sub-summaries (?).
+Another noteworthy technique is the generation of a sketch to structure the final summary along this plan (?).
+Negation remains a long-standing challenge for language models often neglected in computational studies (?, ?).
 4 Datasets
-In this review, we collected both methods and downstream datasets from the selected publications.
+In this review, we collected data generation methods and downstream datasets from the selected publications.
 We only consider datasets used at least five times in our analysis to avoid including resources with little community impact.
-This yields 18 datasets that we group into six categories based on the dialogue subdomains they stem from: daily chat, online chat, meeting, screenplay, customer service, and medical.
-We also create another category for datasets that don’t fit neatly into these subdomain groups.
-These datasets typically condense a short dialogue of about 1k tokens between two real persons into a third-person summary of about 100 tokens.
-More complex scenarios, such as meetings, parliamentary debates, and TV series dialogues, involve around 10k to 20k tokens in the transcript and five or more participants.
-Due to privacy concerns, a significant portion of currently publicly available datasets are reenactments based on real events, e.g., by Amazon Mechanical Turk workers, or established datasets such as SAMSum (?) and AMI (?) are modified to generate new data, e.g., by changing the dialogue transcript to a third-person report (?).
+This setup yields 18 datasets that we group into six categories based on the dialogue subdomains they stem from: daily chat, online chat, meeting, screenplay, customer service, medical, and debates.
+We also create an ‘others’ category for datasets that do not fit neatly into these subdomain groups.
+These datasets typically condense a short dialogue of about 1k tokens between natural persons into a third-person summary of about 100 tokens.
+More complex scenarios, such as meetings, parliamentary debates, and TV series dialogues, involve around 10k to 20k tokens in the transcript and at least four participants.
+Due to privacy concerns, a significant portion of currently publicly available datasets are reenactments based on actual events, e.g., by Amazon Mechanical Turk [^7] https://www.mturk.com/ workers, or established datasets such as SAMSum (?) are modified to generate new data, e.g., by changing the dialogue transcript to a third-person report (?).
 This underscores that the variety of existing datasets is comparably limited, and these datasets may not capture real-life scenarios.
@@
 We did not assess dataset quality or perform a detailed analysis of the inherent challenges, leaving that for future research.
-In Figure 9, we connect the datasets with their respective primary challenge, as noted by their creators, to provide an overview of the distribution of challenges in datasets.
-Figure 9: Matching between established datasets and their primary reported challenge.
+In Table 3, we connect the datasets with their respective primary challenge, as noted by their creators, to provide an overview of the distribution of challenges in datasets.
 4.1 Daily Chat
 DialogSum (?) integrates dialogues from DailyDialog (?), Dream (?), Mutual (?), and English-speaking practice websites, featuring two-speaker interactions across daily-life scenarios like work, leisure, and travel.
-Annotators were given guidelines on writing the summaries, such as length constraints (no longer than 20% of conversation length), maintaining an observer perspective, and explicitly describing important emotions related to events in the summary.
+Annotators were given guidelines on writing summaries, such as length constraints (no longer than 20% of conversation length).
 DialogSum includes 13k dialogues, with inputs of around 1k tokens and summaries of 130 tokens.
@@
 The messages, each written by one person, contain chit-chats, gossip, arranging meetings, discussing politics, and consulting university assignments.
-They typically involve two speakers, with an average of 94 tokens per conversation, varying between three and 30 turns  [^7] A turn is the contribution made by a speaker in the form of a single utterance or a statement..
+They typically involve two speakers, with an average of 94 tokens per conversation, varying between three and 30 turns [^8] A turn is a contribution made by a speaker in the form of a single utterance or a statement..
 A subset of SAMSum has been adapted into formal third-person language (?) to help models transition from informal dialogue to edited text, addressing the linguistic gap between pre-training and downstream task (detailed in the language challenge, Section 3.1).
 Forum (?) consists of random samples of 100 threads from the online discussion forums on Ubuntu and TripAdvisor, with a total of 556 and 916 posts, respectively.
-Two human evaluators created summaries of the discussion threads, resulting in two human written summaries per sample.
-CRD3 (?) consists of dialogues from the ”Critical Role Dungeon and Dragon” show with summaries collected from the Fandom wiki, featuring 159 episodes with an average dialogue length of 2550 turns.
+Two human evaluators created summaries of the discussion threads, resulting in two human-written summaries per sample.
+Crd3 (?) consists of dialogues from the ‘Critical Role Dungeon and Dragon’ show with summaries collected from the Fandom wiki, featuring 159 episodes with an average dialogue length of 2550 turns.
 4.3 Meeting
@@
 ICSI (?) has 59 academic group meetings with computer scientists, linguists, and engineers at the International Computer Science Institute (ICSI) in Berkeley, along with their summaries written by hired annotators.
-The meetings have an average of 819 turns and 13k tokens with six speakers and are about research discussions among students.
+The meetings have an average of 819 turns and 13k tokens with six speakers and are research discussions among students.
 QMSum (?) introduces query-based summarization across diverse meeting domains, compiling 1.8k query-summary pairs from 232 meetings, encompassing product design (AMI), academic discussions (ICSI), and committee deliberations.
@@
 Meetings shorter than half an hour are exceptions, whereas meetings longer than two hours are topic-oriented mini-workshops, also rather occasional.
-The original task here slightly differs as a model is required to produce not an abstractive summary but a set of meeting minutes in bullet points.
+The original task of the dataset differs from abstractive summarization as a model is required to produce not an abstractive summary but a set of meeting minutes in bullet points.
 MeetingBank (?) contains meetings of city councils from six major U.S.
@@
 4.4 Screenplay
-MediaSum (?) contains 463.6k media interview transcripts with brief summaries from National Public Radio (NPR) (?) and CNN, spanning a range of topics.
+MediaSum (?) contains 463.6k media interview transcripts with summaries from National Public Radio (NPR, ?, 2020) and CNN, spanning a range of topics, including politics, news, crime, and economy.
 The summaries are based on NPR’s interview overviews and CNN’s topic descriptions, with the latter segmented at commercial breaks to match topics to corresponding interview segments.
-On average, each transcript in this dataset contains 1.5k words, with summaries of around 11 words, typically involving 7 speakers.
+On average, each transcript in this dataset contains 1.5k words, with summaries of around 11 words, typically involving seven speakers.
 Summscreen (?) consists of 27k instances of TV series transcripts paired with human-written recaps sourced from TV MegaSite and ForeverDreaming.
@@
 TweetSumm (?) contains 1.1k dialogues derived from Twitter customer support exchanges, each with three extractive and three abstractive human-written summaries.
-Originating from the Kaggle Customer Support On Twitter dataset, these real-world interactions span various industries, such as airlines and retail, averaging 10 turns per dialogue and 36 tokens per abstractive summary.
+Originating from the Kaggle Customer Support On Twitter dataset, these real-world interactions span various industries, such as airlines and retail, averaging ten turns per dialogue and 36 tokens per abstractive summary.
 TodSum (?) is a customer service dataset based on MultiWOZ (?), from which they select the five domains: restaurant, hotel, attraction, taxi, and train.
@@
 MTS-Dialog (?) is a collection of 1.7k simulated doctor-patient conversations with corresponding clinical notes serving as summaries, sourced from the public Mtsamples collection (?).
-This dataset encompasses various medical fields, including general medicine, neurology, orthopedics, dermatology, and allergy/immunology, adhering to the SOAP (Subjective, Objective, Assessment, Plan) note format.
+This dataset encompasses various medical fields, including general medicine, neurology, orthopedics, dermatology, and immunology, adhering to the SOAP (Subjective, Objective, Assessment, Plan) note format.
 Dialogues average 18 sentences and 242 words, with summaries typically around 81 words.
@@
 ADSC (?) features sequences of two-party dialogue chains derived from the Internet Argument Corpus (?), focusing on significant social and political topics like gun control, gay marriage, the death penalty, and abortion.
-It includes 45 dialogues, each accompanied by 5 unique human-generated summaries, with each summary approximately 150 words in length.
+It includes 45 dialogues, each accompanied by five unique human-generated summaries, with each summary approximately 150 words in length.
 4.8 Others
@@
 Drawing on SAMSum and DialogSum, Ferranti includes summaries produced by models like BART (?) and Unilm (?).
-Human annotators assess these summaries for accuracy and identify errors, providing a focused tool for improving summary factualness.
-DialSum (?) is based on the VisDial dataset (?), where two participants discuss an image from the MSCOCO dataset (?), which features approximately 120k images.
-Each image thereby has five captions from five different annotators.
+Annotators assess these summaries for accuracy and identify errors, providing a focused tool for improving summary factualness.
+DialSum (?) is based on the VisDial dataset (?), where two participants discuss images from the MSCOCO dataset (?), which features $\sim$120k images.
+Each image has five captions from five different annotators.
+Type Dataset Challenge Descriptive Tags Usage
+Daily Chat DialogSum Speaker daily life scenarios 26
+Online SAMSum Speaker, Salience messaging apps 68
+Chat FORUM Salience threads 5
+CRD3 Language live-streamed show 6
+Meeting AMI Salience, Comprehension staged business meetings 63
+ICSI Salience, Language academic group meetings 21
+QMSum Salience, Language query-based summarization 18
+Elitr Salience technical project meetings in computer science 7
+MeetingBank Structure parliament meetings 7
+Screenplay MediaSum Structure media interview 15
+SummScreen Speaker, Salience TV series transcripts 8
+Customer TweetSumm Language Twitter customer support 7
+Service TODSum Language, Factuality open-domain task-oriented dialogues 6
+Medical MTS-Dialog Comprehension simulated doctor-patient conversations with clinical notes 5
+Debates ADSC Comprehension two-party dialogues on social and political topics 5
+Others ConvoSumm Salience dialogues from comments, emails, and threads 5
+Ferranti Factuality human assessed automatic summaries 5
+DialSum Comprehension two-party discussion on images 10
+Table 3: Matching between established datasets and their primary reported challenge.
+‘Usage’ indicates the number of papers reporting the dataset in their publication.
 4.9 Data Augmentation and Utility
@@
 As discussed earlier, dialogue summarization faces challenges due to the need for adaptability across various domains, structures, and speaker dynamics, which typically would be addressed by training models on diverse datasets (?).
-However, the scarcity and small size of datasets, with utterances sometimes containing just two to ten turns in areas like customer service and medical, restrict model training for more general applications.
+However, datasets’ scarcity and small size, with utterances containing just two to ten turns in areas like customer service and medical, restrict model training for more general applications.
 Creating real-world datasets is costly and may conflict with data security and information communication policies, so artificially generated datasets are explored (?).
 Methods range from simple heuristic-based weak labeling, such as selecting the leading or longest utterance as proxy summary (?), to paraphrasing with updating the summary to maintain coherence (?, ?, ?), to random alterations (e.g., swapping, deletion, dialogue-act-guided insertions) and changing conversation structures (?, ?).
-The most common method is to use language models to create ground truth summaries either directly (?, ?, ?, ?) or after training them on human summarization techniques through few-shot learning (?).
+The most common method is to use language models to create ground truth summaries either directly (?, ?, ?, ?) or after training them on human summarization patterns through few-shot learning (?).
 Techniques to Maximize Dataset Utility.
-As training Transformer-based models requires extensive data to generalize across various topics and dialogue formats (?), researchers have explored ways to use the limited data available more effectively.
+As the training of Transformer-based models requires extensive data to generalize across various topics and dialogue formats (?), researchers have explored ways to use the limited data available more effectively.
 A key strategy involves a prompt-based curriculum learning strategy that gradually increases the degree of prompt perturbation (e.g., word swapping, content cutting) to improve the generalization ability of models (?).
-Another variation is dynamic prompts, which select best-fitting few-shot samples considering dialogue content, size, and speaker number (?).
-Also explored are prompt transfer techniques from related dialogue domains (?) to bolster the usage of dialogue state information.
+Another variation are dynamic prompts, which select best-fitting few-shot samples considering dialogue content, size, and speaker number (?).
+Also explored are prompt transfer techniques from related dialogue domains (?) to bolster the usage of dialogue state information (i.e., data used to represent the underlying intentions and goals within a dialogue).
 Prompts are further used to split inputs into domain-invariant and domain-specific content to enhance model generalization (?, ?).
-Other relevant techniques include freezing most model parameters and training only specific parameters for domain adaption (?, ?, ?).
+Further techniques include freezing most model parameters and training only specific parameters for domain adaption (?, ?, ?).
 5 Evaluation
 Researchers have adopted metrics from related fields and introduced new ones designed to assess the effectiveness of dialogue summarization methods.
-These metrics are used as a proxy for quality characteristics, such as coherence, fluency, factuality, and accuracy.
-In our literature review, we identify 15 metrics used more than twice in 93 papers.
-We categorize these metrics into four groups: count-based (e.g., ROUGE (?)), model-based (e.g., BARTScore (?)), QA-based (e.g., QAEval (?)), and human evaluation.
+These metrics proxy for quality characteristics, such as coherence, fluency, factuality, and accuracy.
+Our literature review identifies 15 metrics used more than twice in 93 papers.
+We categorize these metrics into four groups: count-based (e.g., ROUGE ?, 2004), model-based (e.g., BARTScore ?, 2021), QA-based (e.g., QAEval ?, 2021), and human evaluation.
 The automatic metrics strive for alignment with human judgment.
@@
 Recent analyses (?, ?) conclude that established automatic metrics may work for a superficial understanding of a model’s performance but do not align well with human judgment in discerning error nuances.
-These metrics either do not adequately display error occurrence or cannot reflect the impact on the quality of their scores.
-Hence, a composite metric of count-based, model-based, and QA-based metrics may be required for a more reliable automatic interpretation.
-Section 5.1 details the characteristics typically evaluated by the metrics and Sections 5.2, 5.3, 5.4 and 5.5 further summarizes the 15 established metrics in dialogue summarization.
+These metrics show individual weaknesses where they individually do not adequately reflect occurrence across all error types (e.g., of hallucinated content) or cannot show the impact on the quality in their scores (e.g., when information is omitted).
+Hence, a composite metric of count-based, model-based, and QA-based metrics may be required for a more reliable automatic interpretation, with the individual metrics focusing on specific error types.
+Section 5.1 details the characteristics typically evaluated by metrics and Sections 5.2, 5.3, 5.4 and 5.5 further summarizes the 15 established metrics in dialogue summarization.
 Table 4 provides an overview of the identified metrics and their usage throughout the papers considered for this literature review.
@@
 Low-quality text may contain more redundancy, incoherence, grammatical errors, poor structure, or inappropriate language, while high-quality text would closely align with the reference.
-We identify ten key characteristics (in italic) organized into four overarching groups: accuracy, content, readability, and style.
-Accuracy is the core of a high-quality summary (?), making adherence to the original text’s truth and, therefore, factuality paramount.
-A summary must accurately reflect the facts, events, and details from the source without distortion.
+We identify ten key characteristics discussed in the literature (in italic) and organize them into four overarching groups: accuracy, content, readability, and style.
+Accuracy is the core of a high-quality summary (?), making it crucial to maintain the truth of the original text and ensure factuality.
+A summary must accurately reflect the facts, events, and details from the source without any distortion.
 Content quality is mainly driven by relevance.
@@
 Informativeness goes a step further by selecting crucial information that conveys the depth of the input transcript, enabling readers to grasp the main points without needing to refer back to the original text (?).
-Readability is essential for making summaries accessible and easy to comprehend.
+Readability is essential for making summaries accessible and easy to grasp the key points quickly.
 Coherence ensures that information is presented logically with smooth transitions, making the summary easy to follow (?).
-The structure and organization of a summary further enhances its readability and facilitates information processing, ensuring that it is well-organized and logically constructed (?).
+The structure and organization of a summary further enhances its readability and facilitates information processing, ensuring that it is well-organized and logical (?).
 Conciseness (?) and non-redundancy (?) are also essential as a summary should contain only the essence of the input transcript without unnecessary details.
-Style further contributes to the perceived quality of a text.
-Consistency in perspective, tense, and style throughout the summary contributes to its readability and professional quality (?).
-Fluency focuses on grammar, vocabulary, and sentence structure and aims for a natural, easy-to-read summary that is free from phrasing errors, enhancing the information’s overall clarity and accessibility (?).
+Style pertains to the text’s perceived quality and formal presentation, contributing to its professional polish.
+Consistency in perspective, tense, and stylistic choices throughout the summary contributes to the overall perception of the text’s professionalism (?).
+Fluency focuses on grammar, vocabulary, and sentence structure and aims for a natural, easy-to-read summary free from phrasing errors, enhancing the information’s overall clarity and accessibility (?).
 5.2 Count-Based
 Count-based metrics, including N-gram-based measures like BLEU (?), and ROUGE (?), and statistical measures such as Perplexity (?), are often static and rule-based algorithms that have a long history for evaluating text summarization.
-However, due to their simple approach, they have been criticized for their limitations in capturing overall meaning, fluency, coherence, or factuality in their scores (?).
-BLEU (?) measures the precision of generated summaries against reference texts by examining the overlap of word sequences (typically 1- to 3-word N-grams (?)), emphasizing similarity to the original phrasing but risking a linear relationship to noise (?).
+However, due to their simplicity, they have been criticized for their limitations in capturing overall meaning, fluency, coherence, or factuality (?).
+BLEU (?) measures the precision of generated summaries against reference texts by examining the overlap of word sequences (typically 1- to 3-word N-grams, ?, 2019), emphasizing similarity to the original phrasing but risking a linear relationship to noise (?).
 ROUGE (?) focuses on lexical similarity (?) and aims to capture the extent to which key content from the source is included in the summary, prioritizing content coverage through recall of N-gram overlap, thus attempting to address BLEU’s limitations by emphasizing content inclusion over mere precision.
 A frequently used variation is ROUGE-L, which measures the longest common subsequence between a candidate and a reference.
-METEOR (?) advances BLEU and ROUGE by incorporating both precision and recall, along with synonym matching for semantic analysis, thus enabling to capture the semantic similarity between a candidate and a reference.
+METEOR (?) advances BLEU and ROUGE by incorporating both precision and recall, along with synonym matching for semantic analysis, thus enabling the capturing of the semantic similarity between a candidate and a reference.
 Despite its more balanced approach and sentence-level focus, METEOR, like its predecessors, is prone to noise interference (?).
@@
 With advancements in language models, there is an increasing focus on model-based evaluation metrics due to their higher correlation with human judgment, though count-based metrics remain popular.
-Model-based metrics range from those using handcrafted features and learned embeddings to hybrids combining contextual embeddings with count-based approaches.
-After representing the respective text in a latent space, these metrics consider the similarity between the candidate and reference text, aiming to offer a more nuanced assessment by focusing on semantic similarity, likelihood, and usefulness of summaries.
-However, despite their potential to adapt to changes in language use (e.g., distribution drifts) (?), these metrics can still be error-prone (?), slower than count-based metrics and may not directly measure the specific characteristics outlined in Section 5.1, making it challenging to discern which particular aspect influences their scores.
-BERTScore (?) leverages BERT embeddings (?) to contextually embed the reference and candidate text, calculate the pairwise cosine-similarity on the token level to build a similarity matrix, and greedily pick the highest similarity scores to compute the harmonic mean of precision and recall, thereby capturing semantic nuances (?).
-MoverScore (?) was introduced concurrently to BERTScore following a similar approach, as it also leverages BERT embeddings (?) but instead considers the distance between reference and candidate text, hence employing the Word Mover’s Distance (?), a special case of the Earth Mover’s Distance (?), to measure semantic distance.
-This change in measure allows MoverScore to map semantically related words from one sequence to one word in the other sequence (many-to-one).
+Model-based metrics encompass a wide range of approaches, including those that calculate semantic similarity (e.g., BERTScore, MoverScore), text generation likelihood (e.g., BARTScore), and entailment probability (e.g., FactCC).
+These metrics typically represent text in a latent space using pre-trained embeddings or contextual representations, aiming to provide a more nuanced assessment by focusing on semantic similarity, likelihood of text generation, and factual consistency.
+While these metrics have the potential to adapt to evolving language use (e.g., distribution drifts) (?), they can still be error-prone (?), slower than count-based metrics, and may not directly measure specific characteristics (as outlined in Section 5.1), making it challenging to discern which particular aspect influences their scores.
+BERTScore (?) leverages BERT embeddings (?) to assess textual similarity.
+The metric first contextually embeds the reference and candidate texts, then constructs a similarity matrix through pairwise cosine similarities on the token level.
+The final score is computed by greedily selecting the highest similarity scores and calculating the harmonic mean of precision and recall, enabling the metric to capture semantic nuances beyond surface-level matching (?).
+MoverScore (?) was introduced concurrently to BERTScore following a similar approach, as it also leverages BERT embeddings but instead considers the distance between reference and candidate text, hence employing the Word Mover’s Distanc (?), a special case of the Earth Mover’s Distance (?), to measure semantic distance.
+This change in measure allows MoverScore to map semantically related words from one sequence to one word in another sequence (many-to-one).
 BARTScore (?) evaluates the plausibility of generating a reference text from a given generated text, and vice versa, by calculating the log-likelihood of a sequence that a BART (?) model would typically generate based on the given context.
-This evaluation focuses on assessing both the fluency and the semantic accuracy of the generated content.
+This evaluation focuses on assessing both the fluency and the semantic accuracy.
 BLEURT (?) extends beyond mere embedding comparisons by incorporating a BERT model pre-trained on lexical- and semantic-level supervision signals and fine-tuned on human judgments, enabling it to make detailed assessments of text quality, including coherence and relevance.
@@
 5.4 QA-Based
-The QA-based metrics we identify in the literature concentrate on the factual correctness of a summary using external question-answering systems.
-The effectiveness of these metrics largely depends on the performance of the QA systems, which may not always align with human evaluation standards.
+The QA-based metrics we identify in the literature focus on the factual correctness of a summary by using external question-answering systems.
+These metrics leverage pre-trained transformer-based models to generate questions and assess whether the summary contains the correct answers.
+Their effectiveness largely depends on the quality of the underlying QA systems, which may not always align perfectly with human evaluation standards.
 FEQA (?) generates questions based on the summary content and verifies whether their answers can be found in the source document.
@@
 It expands the QA-based evaluation scope by incorporating various question types and emphasizing the summary’s informativeness.
-QuestEval (?) combines the approaches of FEQA and SummQA, thereby adopting a bidirectional strategy, generating questions from the summary to compare with the source text and vice versa.
+QuestEval (?) combines the approaches of FEQA and SummaQA, thereby adopting a bidirectional strategy, generating questions from the summary to compare with the source text and vice versa.
 This bidirectional evaluation offers a balanced and holistic assessment by considering included and omitted information in the summary.
 Type Category Metric Descriptive Tags Usage
-Count-based N-gram BLEU word overlap, precision, multiple references 21
-ROUGE word overlap, recall, one reference 127
-METEOR word overlap, precision and recall, one reference 3
+Count based N-gram Bleu overlap, precision, multiple references 21
+Rouge overlap, recall, one reference 127
+Meteor overlap, precision and recall, one reference 3
 ChrF++ character-level, F1 score, one reference 2
 Cider consensus-based, multiple references 3
-[…]
-Model-based Hybrid BERTScore token similarity, cosine-similarity, one-to-one 38
-MoverScore token similarity, mover distance, one-to-many 3
-FactCC entailment classifier, scoring based on consistency 9
-[…]
-BARTScore mimics human judgement, promptable 13
-BLANC fill-in-the-blanc task with and without the summary 2
-QA-based FEQA questions based on summary, answered through input 38
+Statistical Perplexity likelihood of word sequences 3
+Model based Hybrid Bertscore token-level, cosine similarity, one-to-one 38
+Moverscore token-level, mover distance, one-to-many 3
+Factcc entailment classifier, scores consistency 9
+Trained Bleurt mimics human judgment 7
+Bartscore mimics human judgment, promptable 13
+Blanc fill-in-the-blank task with and without the summary 2
+QA based Feqa questions based on summary, answered through input 38
 Summaqa questions based on input, answered through summary 3
-QuestEval combination of FEQA and SummQA 3
-Human Evaluation Performance Likert Scale scale, typically 1 (worst) to 5 (best) 6
+Questeval combination of Feqa and Summaqa 3
+Human Evaluation Performance Likert Scale ordinal scale, e.g., 1 (worst) to 5 (best) 6
 Pairwise Comparison pick best sample out of two 9
 Best-Worst Scaling rank a list of samples 5
-[…]
-Cohen’s Kappa measure agreements, categorical data, 2 raters 3
-Fleiss’ Kappa measure agreement, nominal data, 2 raters 2
+Agreement Krippendorff’s Alpha disagreement, different data formats 2
+Cohen’s Kappa agreements, categorical data, 2 raters 3
+Fleiss’ Kappa agreement, nominal data, 2 raters 2
 Table 4: Relevant metrics and evaluation measures employed in dialogue summarization with more than twice reported use.
-’Usage’ states the number of papers reporting the evaluation measure in their publication.
+‘Usage’ states the number of papers reporting the evaluation measure in their publication.
 5.5 Human Evaluation
 Human evaluation is often considered the gold standard for assessing the quality of a summary.
-It is typically performed through crowdsourcing annotators (e.g., Amazon Mechanical Turk) where they label samples.
+It is typically performed through crowdsourcing annotators (e.g., Amazon Mechanical Turk) who label samples.
 We identify established approaches for human evaluations on summary performance and annotator agreement.
@@
 6 Discussion
+Throughout this work, we provide an overview of the current state of challenges, datasets, and evaluation.
+In Section 6.1, we analyze the emerging trends such as LLMs on mitigating the individual challenges, finding that our challenge taxonomy remains up to date.
+Section 6.2 discusses the increasing interest in datasets covering personalized summarization and realistic settings, while Section 6.3 covers recent research on improving evaluation metrics.
 6.1 Remarks on Challenges
@@
 Despite advancements in addressing the challenges, we observe that most of them are still a hurdle for models due to limitations in the Transformer components used in these models, the lack of capabilities in contextualization and few-shot adaption of the encoder-decoder backbone models, and missing mitigation techniques.
-Since 2023, NLP has seen a significant shift with the exploration of LLMs and optimized Transformer architecture components (e.g., ring attention (?) and multi-token generation (?)), with their application to dialogue summarization being explored later (?, ?, ?).
+Since 2023, NLP has seen a significant shift with the exploration of LLMs and optimized Transformer architecture components (e.g., ring attention, ?, 2023, and multi-token generation, ?, 2024), with their application to dialogue summarization being explored later (?, ?, ?).
 Initial studies indicate that LLMs match or exceed the performance of task-specific models like DialogLED (?), even with their limitations (e.g., hallucinations, salience) (?).
@@
 LLMs handle the comprehension challenge but seem to be sub-optimal regarding robustness to language and speakers.
-These early observations state LLMs as a noteworthy alternative to more traditional techniques (e.g., dialogue-style pre-training (?), graph structures to represent speaker structures (?), and special losses tailored to determine salience (?)).
-In the following, we discuss how the hurdles of the currently employed encoder-decoder models (identified in Section 3) align with challenges in NLP and discuss techniques introduced so far to mitigate these challenges, thereby deriving how they can aid dialogue summarization.
+These early observations state LLMs as a noteworthy alternative to more traditional techniques (e.g., dialogue-style pre-training, ?, 2022, graph structures to represent speaker structures, ?, 2022, and special losses tailored to determine salience, ?, 2023).
+Following, we discuss how the hurdles of the currently employed encoder-decoder models (identified in Section 3) align with challenges in NLP and discuss techniques introduced to mitigate these challenges, thereby deriving how they can aid dialogue summarization.
 Bridging the language gap.
@@
 While traditional models could theoretically handle such lengthy contexts, practical limitations arise from the input-length-dependent quadratic computational costs associated with Transformer attention mechanisms (?).
-To address these issues, techniques like sparse attention (?), flash attention (training stage) (?) and ring attention (inference stage) (?), sharing weights across attention heads (?), and conditional computation to reduce the overall computational load (?) are proposed to manage large inputs.
+To address these issues, techniques like sparse attention (?), flash attention during the training stage (?) and ring attention for inference (?), sharing weights across attention heads (?), and conditional computation to reduce the overall computational load (?) are proposed to manage large inputs.
 Generalizability.
@@
 Meanwhile, LLMs build on these techniques to leverage patterns observed during the training stage and boost their generalization capabilities (?).
-However, LLMs may still struggle with bias mitigation and ensuring broad applicability across various contexts, when ”encountered words are not present in the training corpus or in examples (?, ?, ?).
+However, LLMs may still struggle with bias mitigation and ensuring broad applicability across various contexts when encountered words are not present in the training corpus or in examples (?, ?, ?).
 Coreference resolution.
@@
 Avoiding hallucinated content in generated text is a major challenge in NLP tasks using LLMs, especially in culturally sensitive contexts (?, ?).
-Mitigation techniques include confidence penalty regularization (?, ?) to reduce overconfidence and enhance accuracy.
+Mitigation techniques include confidence penalty regularization (?, ?) to reduce overconfidence and enhance accuracy and refinement methods to review and post-process a summary (?).
 These advancements throughout NLP suggest upcoming shifts in the key challenges of dialogue summarization.
   • The language challenge can lose impact if few-shot learning is used to prompt LLMs, making it easier to apply document summarization techniques to dialogues directly without additional finetuning.
-  • We expect progress for the structure challenge with LLMs as the backbone model, but robustness issues may persist due to issues with generalization and processing long texts.
+  • We expect progress for the structure challenge using LLMs, but robustness issues may persist due to issues with generalization and processing long texts.
   • The comprehension challenge, especially in grasping implied meanings, remains largely unsolved and holds significant research potential.
-Research from related NLP fields such as sentiment analysis (?) suggest that current models still struggle to understand implied content.
+Research from related NLP fields such as sentiment analysis (?) suggests that current models still struggle to understand implied content.
 At the same time, research on retrieval-augmented generation (?) could enhance contextualization, helping models better grasp the context of conversations.
   • For the hurdles implied in the speaker challenge, i.e., coreference resolution and dependency parsing, LLMs are explored but have not yet succeeded due to robustness weaknesses (e.g., errors due to uncommon names, overseeing details in long inputs), signaling that further research is required.
-  • The salience challenge currently builds on personalizing summaries and understanding long discussions, which both are gaining attention lately.
-Due to the novelty, the personalization of summaries requires more research on techniques and benchmarking datasets.
+  • The salience challenge builds on personalizing summaries and understanding long discussions, which both are gaining attention lately.
+Due to the novelty, the personalized summaries require more research on techniques and benchmarking datasets.
   • The ongoing issue of hallucination in language models continues to pose a challenge in ensuring the factuality in generated summaries.
@@
 Personalized Summarization.
-In contrast to the typically general summaries, personalized summaries tailored to the users’ needs are becoming increasingly popular (?, ?).
+In contrast to the typically general summaries, personalized summaries tailored to the users’ needs are becoming increasingly popular (?, ?, ?).
 Recent advancements have explored cross-attention and decoder self-attention to enhance role-specific information capturing (?).
@@
 This set of metrics thereby mimics human judgment and assesses common aspects (e.g., fluency, coverage, coherence) with continuous (?), Likert scale (?, ?), or probability scores (?).
-LLM-based metrics offer a promising direction for dialogue summarization evaluation due to their customizability, though this area remains largely unexplored.
+LLM-based metrics offer a promising direction for dialogue summarization evaluation due to their customizability (?), though this area remains largely unexplored.
 7 Final Considerations
-7.1 Limitations of Evidence
-Given our selection of papers on dialogue summarization, we found limitations in current works that may bias our findings.
-First, the crawled research studies in dialogue summarization predominantly focus on English-language dialogues, which renders their performance in multilingual or cross-cultural contexts unknown.
-While we frequently came across Chinese datasets, other languages are underrepresented.
-As a consequence, we focus our work on English dialogue summarization.
-Second, most datasets stem from specific sectors such as customer service or business meetings, which may not reflect the variety of dialogue types found in different environments.
-Lastly, the lack of detailed human evaluation reporting across research works, including inter-annotator agreement scores, further complicates the assessment of new techniques’ usefulness.
-7.2 Limitations of Review Process
-In our systematic review of dialogue summarization, we encountered methodological challenges that we had to circumvent.
-First, we focus our search on studies published in English and primarily involving English datasets.
-While this is a limiting factor, most research on the dialogue summarization task considers only English dialogues, making this a feasible approximation.
-However, we encourage others to assess the same issues we point out for other languages.
-Second, despite considering multiple databases, the timing and methods of our search may have missed studies, thus not fully capturing the dialogue summarization field.
-To mitigate this limitation, we updated our crawl several times throughout writing this article and retrieved works from two databases.
-We considered both peer-reviewed and non-peer-reviewed papers which risks including less credible studies or misinformation.
-To mitigate the potential weaknesses of non-peer-reviewed works, we use the ranking from semantic scholar and DBLP, which should account for the paper’s quality and evaluate the studies’ quality based on adherence to established reporting standards of top-tier conferences.
-7.3 Conclusion
 In this article, we reviewed works on Transformer-based abstractive dialogue summarization.
@@
 We conclude that while challenges like language may become less relevant, others, such as comprehension and factuality, still require more exploration.
-We recommend that future work should further discuss and adapt this taxonomy as new challenges, techniques, datasets, and evaluation measures emerge, and as current challenges may be solved naturally through technological innovation.
-The considered literature can be found in the accompanying repository  [^8] https://anonymous.4open.science/r/LitRev-DialogueSum/ which will be regularly updated.
+We recommend that future work should further discuss and adapt this taxonomy as new challenges, techniques, datasets, and evaluation measures emerge and as current challenges may be solved naturally through technological innovation.
+The considered literature can be found in the accompanying repository [^9] https://github.com/FKIRSTE/LitRev-DialogueSum which will be regularly updated.
+In the following, we discuss the discuss limitations of evidence (Section 7.1) and in our review process (Section 7.2).
+7.1 Limitations of Evidence
+Given our selection of papers on dialogue summarization, we found limitations in current works that may bias our findings.
+First, the crawled research studies in dialogue summarization predominantly focus on English-language dialogues, which renders their performance in multilingual or cross-cultural contexts unknown.
+While we frequently came across Chinese datasets, other languages are underrepresented.
+As a consequence, we focus our work on English dialogue summarization.
+Second, most datasets stem from specific sectors, such as customer service or business meetings, which may not reflect the various dialogue types in different environments.
+Lastly, the lack of detailed human evaluation reporting across research works, including inter-annotator agreement scores, further complicates the assessment of new techniques’ usefulness.
+7.2 Limitations of Review Process
+In our systematic review of dialogue summarization, we encountered methodological challenges that we had to mitigate.
+First, we focused our search on studies published in English and mainly using English datasets.
+While this is a limiting factor, most research on dialogue summarization considers only English dialogues, making this a feasible approximation.
+However, we encourage others to assess the issues we point out for other languages.
+Second, despite considering multiple databases, the timing and methods of our search may have missed studies, thus not fully capturing the dialogue summarization field.
+To mitigate this limitation, we updated our crawl several times throughout writing this article and retrieved works from two databases.
+We considered both peer-reviewed and non-peer-reviewed papers, which risks including less credible studies or misinformation.
+To mitigate the potential weaknesses of non-peer-reviewed works, we use the ranking from semantic scholar and DBLP, which should account for the paper’s quality and evaluate the studies’ quality based on adherence to established reporting standards of top-tier conferences.
 Acknowledgments
-This work was supported by the Lower Saxony Ministry of Science and Culture and the VW Foundation. Frederic Kirstein was supported by the Mercedes-Benz AG Research and Development.
-Glossary
-LLM
-Large Language Model
-NLP
-Natural Language Processing
+This work was supported by the Lower Saxony Ministry of Science and Culture and the VW Foundation. The Mercedes-Benz AG Research and Development supported Frederic Kirstein.